### v1

# OmniDocBench: Benchmarking Diverse PDF Document Parsing with Comprehensive Annotations

Abstract
Document content extraction is crucial in computer vision, especially for meeting the high-quality data needs of large language models (LLMs) and retrieval-augmented generation (RAG) technologies. However, current document parsing methods suffer from significant limitations in terms of diversity and comprehensive evaluation. To address these challenges, we introduce OmniDocBench, a novel multi-source benchmark designed to advance automated document content extraction. OmniDocBench includes a meticulously curated and annotated high-quality evaluation dataset comprising nine diverse document types, such as academic papers, textbooks, slides, among others. Our benchmark provides a flexible and comprehensive evaluation framework with 19 layout category labels and 14 attribute labels, enabling multi-level assessments across entire datasets, individual modules, or specific data types. Using OmniDocBench, we perform an exhaustive comparative analysis of existing modular pipelines and multimodal end-to-end methods, highlighting their limitations in handling document diversity and ensuring fair evaluation. OmniDocBench establishes a robust, diverse, and fair evaluation standard for the document content extraction field, offering crucial insights for future advancements and fostering the development of document parsing technologies. The codes and dataset is available in https://github.com/opendatalab/OmniDocBench.
footnotetext: The authors contributed equally.footnotetext: Project lead.footnotetext: Corresponding author (heconghui@pjlab.org.cn).
1 Introduction
Figure 1: OmniDocBench Data Diversity. It contains 9 PDF page types, along with Layout Annotations and Recognition Annotations. Furthermore, there are 5 Page Attributes, 3 Text Attributes, and 6 Table Attributes.
Document parsing is a foundational task in computer vision, focused on accurately extracting content from documents [18, 36, 39, 41, 45]. High-quality document content extraction typically involves the integration of multiple algorithmic modules. Layout detection algorithms identify different content areas on a page, OCR technology converts images of text regions into text, while formula and table recognition models identify specific regions and transform them into corresponding source code. These modules and reading order algorithms form a comprehensive process of converting documents into machine-readable formats.
With large models increasingly requiring high-quality data, the importance of document content extraction has become more pronounced. Although vast amounts of data are available online for training, knowledge-rich document data is relatively scarce. Documents such as academic papers and technical reports contain rich structured information that can significantly enhance the knowledge depth of large models. Moreover, the development of retrieval-augmented generation (RAG) [21, 10] technology relies on extracting accurate information from documents to improve the quality and relevance of generated content. Consequently, research in document content extraction has intensified, leading to a series of pipeline-based high-quality document extraction algorithms [36] and the emergence of end-to-end multimodal large model solutions [5, 40, 27, 39, 3, 6, 42]. These methods have significantly improved document content parsing quality, providing robust support for the needs of large models and RAG technology.
In analyzing current module-based pipeline and multimodal end-to-end methods, we identified several limitations. For instance, methods like Marker and MinerU, which are mainstream pipeline methods, primarily evaluate individual modules on academic paper data, lacking document diversity and comprehensive evaluation results. Although MinerU considers the generalization of diverse data, it only demonstrates this through a single model and visualization results, lacking overall end-to-end evaluation. Multimodal large model methods [5, 40, 27, 39, 3], while easier to use than pipeline methods, lack performance validation on diverse documents, and some evaluation metrics are inadequate. Additionally, these methods often use data similar to their training set distribution for comparison, resulting in unfair evaluations. Overall, current document content extraction faces the following challenges:
  • Limited document types. Current evaluations mostly focus on a single type of academic paper, while real-world scenarios include textbooks, exam papers, financial reports, newspapers, magazines, and other document types.
  • Monotonous evaluation dimensions. Pipeline-based methods typically evaluate specific algorithmic modules, such as OCR, layout detection, or formula recognition, while the overall quality of parsing results requires comprehensive metrics.
  • Inadequate evaluation metrics. Multimodal large model approaches attempt to evaluate document parsing quality across multiple dimensions, such as dividing document content into text, formulas, tables, etc. However, these models commonly employ evaluation metrics such as BLEU scores or Edit distances, which fail to accurately and fairly assess parsing effectiveness when dealing with markup languages like LaTeX or HTML that allow diverse syntactic expressions.
Building a diverse, comprehensive, and accurate evaluation system poses significant challenges, requiring diverse and high-quality data annotation and reasonable evaluation metrics. While READOC extends the evaluation scope to include GitHub README files based on Nougat, there remains a substantial gap in real-world diversity, and the evaluation dimensions lack consideration of attributes. In contrast, this paper proposes a document content extraction benchmark, OmniDocBench, characterized by diverse types, detailed annotations, and reasonable evaluation (Figure 1). The specific contributions are as follows:
  • High-quality, diverse evaluation set: Through automated annotation, manual verification, and expert review, we construct a comprehensive, detailed, high-quality OmniDocBench evaluation set, encompassing nine types of diverse document pages, including papers, textbooks, exam questions, and research reports.
  • Flexible and comprehensive evaluation dimension support: The OmniDocBench validation set covers 19 layout category labels and 14 attribute labels. To facilitate user evaluation from an overall, single module, or different data types, we provide end-to-end evaluation, single algorithm module evaluation, and attribute-based evaluation, covering various evaluation needs.
  • Comprehensive evaluation of mainstream methods: Based on OmniDocBench, we conduct a comprehensive evaluation of current mainstream modular pipeline and end-to-end large model methods, providing a fairer assessment of existing methods and summarizing the shortcomings of current document parsing methods, thereby guiding further development in document parsing.
2 Related Work
Benchmark Document Categories Annotaion Type Single-Module Eval End-to-End Eval
BBox Text Table Order Formula OCR DLA TR MFR OCR ROD TR MFR
Single-Module Eval Benchmark
Robust Reading [19] 1
PubLayNet [43], DocBank [24], DocLayNet [31], M6Doc [7] 1, 1, 5, 6
PubTabNet [47],TableX [9], TableBank [23] 1, 1, 1
Im2Latex-100K [8],UniMER-Test [34] 1
End-to-end Eval Benchmarks
Nougat [5] 1
Fox [27] 2
GOT OCR 2.0 [39] 2
READoc [26] 2
OmniDocBench 9
Table 1: A comparison between OmniDocBench and existing DCE benchmarks. OCR: Optical Character Recognition; DLA: Document Layout Analysis; MFR: Math Formula Recognition; TR: Table Recognition; ROD: Reading Order Detection
2.1 Traditional Document Content Extraction
Document Content extraction remains a challenging task, and there is yet to emerge a unified benchmark tailored for real-world scenarios. Traditional algorithms typically employ multiple expert modules to handle different extraction subtasks, such as document layout detection [16, 46, 32, 11], optical character recognition (OCR) [33, 22, 28, 14, 37], formula recognition [44, 25, 4, 34], and table recognition [15, 17, 22].
While expert models of these subtasks are advancing rapidly, recent work such as Mineru [36] attempts to concatenate multiple expert modules into a pipeline and provides a high-precision open-source solution for document content extraction. READOC [26] also unifies heterogeneous evaluation methods from the perspective of Document Structure Extraction, breaking down texts, images, formulas, tables, and other dimensions for evaluation, thus offering a solution-oriented towards real-world scenarios for DSE tasks. However, due to the complexity of Document data sources and the intricacies of PDF document information, previous efforts still fall short in terms of data diversity, failing to cover the categories users encounter in practical applications. Similarly, there is an issue with the explainability of document parsing.
2.2 VLM-based Document Content Extraction
The emergence of Vision-Language Models (VLMs) [38, 6, 1, 12] has revolutionized the field of document content extraction. These models leverage multi-modality capability to achieve remarkable performance in document understanding tasks. Document extraction tools powered by VLMs excel at comprehending both visual layouts and textual content, effectively handling complex document structures while capturing rich contextual information. Representative works such as Nougat [5], Vary [40], Fox [27], and GOT [39], along with recent advances [13, 29], demonstrate significant progress in automated document parsing and comprehension. Despite these advances, the field lacks a standardized and unified benchmark for evaluating VLM-based document extraction task. This absence has hindered objective assessment of PDF document processing capabilities and impeded fair comparison across different approaches. To address this limitation, we present OmniDocBench, a comprehensive end-to-end benchmark designed specifically for evaluating VLM-based document parsing in real-world scenarios.
2.3 Benchmark for Document Content Extraction
An end-to-end benchmark for PDFs can intuitively reflect the effectiveness of PDF extraction tools, which is crucial for their iteration and selection. However, current benchmarks predominantly focus on module-level evaluations; we have listed related benchmarks in Table 1. Additionally, while there are existing end-to-end benchmarks, they lack detailed annotation rules and suffer from insufficient diversity, as well as unreasonable metrics for formula and table evaluations. For example, READOC [26] covers only two types of sources—arXiv and GitHub—and uses EDS [20] and TEDS [47] to compute metrics for formulas and tables, which may lead to inaccuracies CDM [35]. Therefore, there is a need for a more finely annotated, diverse, and reasonably evaluated end-to-end benchmark.
3 OmniDocBench Dataset
Constructing a diverse and comprehensive document parsing benchmark with precise annotations is a formidable challenge. As illustrated in Figure 2, we have designed a systematic and professional annotation framework for OmniDocBench, encompassing data acquisition, intelligent pre-annotation, and manual refinement. This ensures that OmniDocBench possesses the following key attributes:
  • Page Diversity. We sourced document pages from a variety of origins to ensure a wide range of document types.
  • Comprehensive Annotation. We meticulously annotated all elements on the pages, including bounding boxes, specific content, and various potential attributes.
  • Annotation Accuracy. By integrating semi-automated annotation processes, annotator corrections, and expert quality checks, we ensure the reliability of all annotations.
The following sections detail the data acquisition process, the annotation methodology, and a statistical analysis of the final annotated dataset.
3.1 Data Acquisition
During the data acquisition phase, we sourced document pages from diverse origins and used clustering algorithms to initially select visually diverse pages, followed by manual annotation of page attributes to finalize the OmniDocBench pages. Specifically, we collected 200,000 initial PDF documents from Common Crawl, Google, Baidu search engines, and internal data. Subsequently, we extracted visual features from these document pages using ResNet-50 and performed clustering using Faiss  [^1] https://github.com/facebookresearch/faiss, sampling 6,000 visually diverse pages from 10 cluster centers. Finally, annotators provided page-level attribute annotations, including page type, layout type, and language type, and further balanced the selection to 981 samples for the final dataset. The OmniDocBench dataset includes pages with 9 types of pages, multiple layout categories, and various attribute annotations, covering a wide range of real-world scenarios.
Figure 2: Overview of the OmniDocBench dataset construction.
3.2 Data Annotation
To ensure the comprehensiveness of OmniDocBench’s annotations, we conducted detailed annotations for layout detection and content recognition.
3.2.1 Annotation Types
Layout Detection Annotations: Unlike typical layout detection tasks, OmniDocBench includes four comprehensive types of annotations: (1) Layout Bounding Box Annotations: Locating information for 19 types of regions such as titles, text paragraphs, tables, and images. (2) Layout Attribute Annotations: Detailed attribute annotations for detected boxes, including 3 text box attributes, 6 table attributes, and 2 formula attributes. (3) Reading Order Annotations: Annotating the reading sequence of detected boxes. (4) Affiliation Annotations: For images, tables, formulas, and code blocks, we annotate captions and titles to distinguish them from main text. Similarly, for cross-page paragraphs, we annotate affiliation relationships.
Content Recognition Annotations: Based on the format of the content area, we conduct the following three types of area annotations: (1) Text Annotations: Pure text annotations for titles, text paragraphs, and other plain text content. (2) Formula Annotations: LaTeX format annotations for inline formulas, display formulas, and subscripts. (3) Table Annotations: Providing both HTML and LaTeX annotations for table data.
3.2.2 Annotation Process
For these annotation tasks on diverse pages, we design a standardized process to ensure quality and efficiency, comprising intelligent pre-annotation, annotator correction, and expert quality inspection.
Intelligent Pre-Annotation. Manually annotating entire documents is time-consuming and costly. To enhance efficiency, we employ state-of-the-art detection and recognition models for pre-annotation of layout detection and content recognition. Specifically, we use fine-tuned LayoutLMv3 [16] for layout detection annotations and PaddleOCR [22], UniMERNet [34], and GPT-4o [2] for text, formula, and table annotations, respectively.
Annotator Correction. After layout detection phase, annotators refine the detection boxes and enhance annotations with reading order and affiliation details. Each character is verified to ensure accuracy in content recognition. For complex annotations of tables and formulas, requiring LaTeX and HTML formats, annotators use tools like Tables Generator  [^2] https://www.tablesgenerator.com/ and latexlive  [^3] https://www.latexlive.com/ for verification and correction.
Expert Quality Inspection. Despite thorough annotator corrections, the complexity of formulas and tables may result in residual issues. To address these, we use CDM’s rendering techniques to identify unrenderable elements. These are then reviewed and corrected by three researchers to ensure accuracy and fidelity in the final annotations.
3.3 Dataset Statistics
Page Diversity. OmniDocBench comprises a total of 981 PDF pages across 9 distinct types. Each page is annotated with global attributes, including text language, column layout type, and indicators for blurred scans, watermarks, and colored backgrounds.
Annotation Diversity: OmniDocBench contains over 10,0000 annotations for page detection and recognition: (1) More than 20,000 block-level annotations across 15 categories, including over 9,000 text paragraphs, 989 image boxes, 428 table boxes, and so on. All document components except headers, footers, and page notes are labeled with reading order information, totaling over 16,000 annotations. (2) The dataset also includes more than 80,000 span-level annotations across four categories, with 4,000 inter-line formulas and footnote markers represented in LaTeX format, while the remaining annotations are in text format.
Annotation Attribute Diversity: (1) Text Attributes: All block-level annotations, except for tables and images, include text attribute tags. In addition to standard Chinese and English text, there are over 2,000 blocks with complex backgrounds and 146 with rotated text. (2) Table Attributes: Besides standard Chinese and English tables, there are 142 with complex backgrounds, 81 containing formulas, 150 with merged cells, and 7 vertical tables.
4 OmniDocBench Evaluation Methodology
To provide a fair and comprehensive evaluation for various models, we proposed an end-to-end evaluation pipeline consisting of several modules, including extraction, matching algorithm, and metric calculation, as shown in Figure 3. It ensures that OmniDocBench automatically performs unified evaluation on end-to-end DCE tasks, thereby producing reliable and effective evaluation results.
Figure 3: OmniDocBench Evaluation Pipeline.
4.1 Extraction
Preprocessing: The model-generated markdown text should be preprocessed, which includes removing images, eliminating markdown tags at the beginning of the document, and standardizing the number of repeated characters.
Special Component Extraction: Extraction is primarily carried out using regular expression matching. To ensure that the extraction of content does not interfere with each other, it is necessary to follow a specific order. The extraction sequence is as follows: LaTeX tables, HTML tables, display formulas, markdown tables (which are then converted into HTML format), and code blocks.
Pure Text Extraction: After extracting special components, the remaining content is considered pure text. Paragraphs are separated by double line breaks, allowing them to participate in subsequent matching processes, thus aligning with reading order annotation units in the GTs. If no double line break exists, single line breaks are used for paragraph separation. Additionally, previously extracted code blocks are merged into the text category for processing.
Inline Formula Format Converting: We standardized inline formulas within paragraphs to Unicode format. This was necessary because different models produce inconsistent outputs for inline formulas. For formulas originally written in Unicode, it is hard to extract them using regular expressions. Therefore, to ensure a fair comparison, we do not extract inline formulas for separate evaluation. Instead, we include them in their Unicode format alongside the text paragraphs for evaluation.
Reading Order Extraction: Upon completion of the extraction, the start and end positions of the extracted content in the original markdown are recorded for subsequent reading order calculation.
[…]
EN ZH EN ZH EN ZH EN ZH EN ZH EN ZH EN ZH
Pipeline Tools MinerU 0.058 0.211 0.278 0.577 66.9 49.5 79.4 62.7 0.305 0.461 0.079 0.288 0.180 0.384
Marker 0.141 0.303 0.667 0.868 18.4 12.7 54.0 45.8 0.718 0.763 0.138 0.306 0.416 0.560
Mathpix 0.101 0.358 0.306 0.454 71.4 72.7 77.9 68.2 0.322 0.416 0.105 0.275 0.209 0.376
Expert VLMs GOT-OCR 0.187 0.315 0.360 0.528 81.8 51.4 53.5 48.0 0.521 0.594 0.141 0.28 0.302 0.429
Nougat 0.365 0.998 0.488 0.941 17.4 16.9 40.3 0.0 0.622 1.000 0.382 0.954 0.464 0.973
General VLMs GPT4o 0.144 0.409 0.425 0.606 76.4 48.2 72.8 63.7 0.363 0.474 0.128 0.251 0.265 0.435
Qwen2-VL 0.252 0.251 0.468 0.572 54.9 60.9 59.9 66.8 0.591 0.587 0.255 0.223 0.392 0.408
InternVL2 0.353 0.290 0.543 0.701 69.8 49.6 63.8 61.1 0.616 0.638 0.317 0.228 0.457 0.464
Table 2: Comprehensive evaluation of document parsing algorithms on OmniDocBench: performance metrics for text, formula, table, and reading order extraction, with overall scores derived from ground truth comparisons.
Model Type Models Book Slides Financial Report Textbook Exam Paper Magazine Academic Papers Notes Newspaper Average
Pipeline Tools MinerU 0.044 0.124 0.033 0.102 0.159 0.072 0.025 0.984 0.148 0.188
Marker 0.188 0.327 0.087 0.292 0.423 0.134 0.102 0.470 0.270 0.255
Mathpix 0.131 0.168 0.202 0.199 0.278 0.138 0.091 0.631 0.648 0.276
Expert VLMs GOT-OCR 0.105 0.222 0.067 0.132 0.204 0.198 0.179 0.388 0.771 0.252
Nougat 0.734 0.958 1.000 0.820 0.930 0.83 0.214 0.991 0.871 0.816
General VLMs GPT4o 0.157 0.163 0.348 0.187 0.281 0.173 0.146 0.607 0.751 0.313
Qwen2-VL 0.094 0.08 0.145 0.148 0.219 0.065 0.315 0.298 0.79 0.239
InternVL2 0.216 0.098 0.162 0.184 0.247 0.150 0.419 0.226 0.903 0.289
Table 3: End-to-end text recognition performance on OmniDocBench: evaluation using edit distance across 9 PDF page types.
Models Fuzzy Water Color Mean Variance
Pipeline Tools
MinerU 0.15 0.151 0.107 0.136 0.0004
Marker 0.286 0.436 0.290 0.337 0.0049
Mathpix 0.294 0.290 0.182 0.255 0.0027
Expert VLMs
GOT-OCR 0.175 0.190 0.186 0.184 0.0000
Nougat 0.934 0.915 0.873 0.907 0.0006
General VLMs
GPT4o 0.263 0.195 0.184 0.214 0.0012
Qwen2-VL 0.101 0.157 0.114 0.124 0.0006
InternVL2 0.120 0.197 0.155 0.157 0.0010
Table 4: End-to-end text recognition on OmniDocBench: evaluation under various page attributes using the edit distance metric. Columns represent: Fuzzy (Fuzzy scan), Water (Watermark), Color (Colorful background).
Models Single Double Three Complex Mean Variance
Pipeline Tools
MinerU 0.311 0.101 0.117 0.376 0.226 0.0143
Marker 0.231 0.251 0.309 0.378 0.292 0.0033
Mathpix 0.189 0.175 0.225 0.413 0.250 0.0091
Expert VLMs
GOT-OCR 0.163 0.145 0.257 0.468 0.258 0.0165
Nougat 0.852 0.601 0.662 0.873 0.747 0.0139
General VLMs
GPT4o 0.109 0.204 0.254 0.426 0.248 0.0132
Qwen2-VL 0.098 0.248 0.517 0.429 0.323 0.0263
InternVL2 0.082 0.312 0.682 0.444 0.380 0.0472
Table 5: End-to-end reading order evaluation on OmniDocBench: results across different column layout types using Normalized Edit Distance.
4.2 Matching Algorithm
To avoid the impact of paragraph splitting on the final results, we proposed a method, Adjacency Search Match, that merges and splits paragraphs in both GTs and Preds to achieve the best possible match. The specific strategy involves: i) Calculate a metrix of Normalized Edit Distance between GTs and Preds. If the similarity between a Pred and a GT exceeds a specific threshold, they are considered a successful match. ii) For the rest, we apply fuzzy matching to determine whether one string is a subset of another string. If so, we further apply the truncation and merging algorithm which would try to merge adjacent paragraph. This process would continue to merge more paragraph until the Normalized Edit Distance starts to decrease. After this process, the best match will be found for GTs and Preds.
4.3 Metric Calculation
Ignore Handling: We implement an ignore logic for certain components in PDF page content, meaning they participate in matching but are excluded from metric calculations. This is mainly because of inconsistent output standards among models, which should not affect the validation results. For fairness, we ignore: (1) Headers, footers, page numbers, and page footnotes, which are handled inconsistently by different models. (2) Captions for figures, tables, and footnotes often have uncertain placement, complicating reading order. Additionally, some models embed table captions in HTML or LaTeX tables, while others treat them as plain text.
Metric: Different calculation methods are used for various document components: (1) Pure Text: We calculate Normalized Edit Distance, averaging these metrics at the sample level to obtain the final scores. (2) Tables: All tables are converted to HTML format before calculating the TEDS metric and Normalized Edit Distance. (3) Formulas: Formulas are currently evaluated using the CDM [35], Normalized Edit Distance, and BLEU. We did not convert interline formulas into Unicode because Unicode cannot represent certain complex formulas, such as matrices. (4) Reading Order: Reading order use the Normalized Edit Distance as metric. It only involves text components, where tables, images, and ignored components do not participate in the final reading order calculation.
5 Benchmarks
Model Book Slides Research Report Textbook Exam Paper Magazine Academic Literature Notes Newspaper Average mAP
DiT-L 43.44 13.72 45.85 15.45 3.40 29.23 66.13 0.21 23.65 26.90
LayoutLMv3 42.12 13.63 43.22 21.00 5.48 31.81 64.66 0.80 30.84 28.84
DOCX-Chain 30.86 11.71 39.62 19.23 10.67 23.00 41.60 1.80 16.96 21.27
DocLayout-YOLO 43.71 48.71 72.83 42.67 35.40 51.44 66.84 9.54 57.54 48.71
Table 6: Component-level layout detection evaluation on OmniDocBench layout subset: mAP results by PDF page type.
5.1 Component-specific Evaluation Results
Model Type Model Language Table Frame Type Special Situation Overall
EN ZH Mixed Full Omission Three Zero Merge Cell(+/-) Formula(+/-) Colorful (+/-) Rotate(+/-)
OCR-based Models PaddleOCR 76.8 71.8 80.1 67.9 74.3 81.1 74.5 70.6/75.2 71.3/74.1 72.7/74.0 23.3/74.6 73.6
RapidTable 80.0 83.2 91.2 83.0 79.7 83.4 78.4 77.1/85.4 76.7/83.9 77.6/84.9 25.2/83.7 82.5
Expert VLMs StructEqTable 72.0 72.6 81.7 68.8 64.3 80.7 85.0 65.1/76.8 69.4/73.5 66.8/75.7 44.1/73.3 72.7
GOT-OCR 72.2 75.5 85.4 73.1 72.7 78.2 75.7 65.0/80.2 64.3/77.3 70.8/76.9 8.5/76.3 74.9
General VLMs Qwen2-VL-7B 70.2 70.7 82.4 70.2 62.8 74.5 80.3 60.8/76.5 63.8/72.6 71.4/70.8 20.0/72.1 71.0
InternVL2-8B 70.9 71.5 77.4 69.5 69.2 74.8 75.8 58.7/78.4 62.4/73.6 68.2/73.1 20.4/72.6 71.5
Table 7: Component-level Table Recognition evaluation on OmniDocBench table subset. (+/-) means with/without special situation.
Model Type Model Language Text background Text Rotate
EN ZH Mixed White Single Multi Normal Rotate90 Rotate270 Horizontal
Expert Vision Models PaddleOCR 0.071 0.055 0.118 0.060 0.038 0.085 0.060 0.015 0.285 0.021
Tesseract OCR 0.179 0.553 0.553 0.453 0.463 0.394 0.448 0.369 0.979 0.982
Surya 0.057 0.123 0.164 0.093 0.186 0.235 0.104 0.634 0.767 0.255
GOT-OCR 0.041 0.112 0.135 0.092 0.052 0.155 0.091 0.562 0.966 0.097
Mathpix 0.033 0.240 0.261 0.185 0.121 0.166 0.180 0.038 0.185 0.638
Vision Language Models Qwen2-VL 0.072 0.274 0.286 0.234 0.155 0.148 0.223 0.273 0.721 0.067
InternVL2 0.074 0.155 0.242 0.113 0.352 0.269 0.132 0.610 0.907 0.595
GPT4o 0.020 0.224 0.125 0.167 0.140 0.220 0.168 0.115 0.718 0.132
Table 8: Component-level evaluation on OmniDocBench OCR subset: results grouped by text attributes using the edit distance metric.
Models CDM ExpRate@CDM BLEU Norm Edit
GOT-OCR 74.1 28.0 55.07 0.290
Mathpix 86.6 2.8 66.56 0.322
Pix2Tex 73.9 39.5 46.00 0.337
UniMERNet-B 85.0 60.2 60.84 0.238
GPT4o 86.8 65.5 45.17 0.282
InternVL2 67.4 54.5 47.63 0.308
Qwen2-VL 83.8 55.4 53.71 0.285
Table 9: Component-level formula recognition evaluation on OmniDocBench formula subset.
The OmniDocBench dataset features comprehensive and precise annotations, allowing for a fair and rigorous comparison of various document content extraction algorithms in real-world scenarios. Based on the distinct characteristics of these algorithms, we categorize document content extraction methods into three main classes:
Pipeline Tools. These methods integrate layout detection and various content recognition tasks (such as OCR, table recognition, and formula recognition) into a document parsing pipeline for content extraction. Prominent examples include MinerU [36], Marker [30], and Mathpix  [^4] https://mathpix.com/.
Expert VLMs. These are large multimodal models specifically trained for document parsing tasks. Representative models include GOT-OCR2.0 [39] and Nougat [5].
General VLMs. These are general-purpose large multimodal models inherently capable of document parsing. Leading models in this category include GPT-4o [2], Qwen2-VL [38], and InternVL2 [6].
5.2 End-to-End Evaluation Results
Utilizing the OmniDocBench dataset and our evaluation framework, we conducted end-to-end assessments of mainstream document parsing methods, evaluating their performance from input PDF images to the resultant document parsing outputs.
Overall Evaluation Results. As illustrated in Table 2, pipeline tools specifically designed for document parsing, demonstrate superior performance across the board. MinerU and Mathpix achieved the best results for English and Chinese pages, respectively. In contrast, even the best general-purpose Vision Language Models (VLMs), GPT-4o, exhibits a performance gap compared to these specialized models, especially in Chinese. This trend is evident across sub-tasks like text recognition, formula recognition, and table recognition, where methods tailored for document parsing consistently outperform others. This advantage is largely due to the fine-tuning of these models on large datasets specific to document parsing tasks.
Performance Across Diverse Page Types. To gain deeper insights into model performance on diverse document types, we evaluated text recognition tasks across different page types. As shown in Table 3, an intriguing finding emerged: For commonly used data, such as academic papers and financial reports, pipeline tools perform well. However, for more specialized data like slides and handwritten notes, general VLMs demonstrate stronger generalization. The reason is clear: Pipeline tools and expert VLMs are relatively more constrained by the range of training data, whereas general VLMs having been trained on a wide variety of samples, maintained excellent recognition performance even in traditionally challenging long-tail scenarios, underscoring the value of VLMs.
Performance on Pages with Specific Attributes. For documents in OmniDocBench with attributes such as fuzzy scans, watermarks, and colorful backgrounds, our evaluation results are presented in Table 4. In these scenarios, the VLMs InternVL2 and Qwen2-VL exhibit the strongest resistance to interference, achieving the best accuracy and robustness. MinerU also performs commendably.
Performance on Different Column Layout Types. OmniDocBench annotates page attributes such as column layout type, which is crucial for analyzing model performance in reading order. As depicted in Table 5, all models experience a noticeable decline in reading order accuracy when dealing with complex layouts. MinerU and Mathpix excels in reading order across various column layouts, demonstrating robust performance across different page types.
From these end-to-end evaluations, it is evident that pipeline tools like MinerU and Mathpix, specifically designed for document parsing, achieve the best overall performance. However, in terms of versatility and scalability, VLMs offer a distinct advantage over pipeline tools. Fine-tuning a general large model like Qwen2-VL with specialized data could yield models even more adept at document parsing, indicating a promising direction for future research in multimodal approaches.
The OmniDocBench dataset provides comprehensive annotations for document parsing, including layout detection, text boxes and content, formula boxes and content, and table boxes and content. These detailed annotations enable the evaluation of current state-of-the-art (SOTA) methods across various document types, allowing us to analyze their performance in diverse scenarios. Additionally, these results can be used to assemble enhanced pipeline tools for document parsing tasks.
5.3 Single Algorithm Evaluation Results
Layout Detection Results. Layout detection is the first step in document parsing using pipeline tools. A robust layout detection algorithm should perform well across a variety of document types. Table 6 presents an evaluation of leading layout detection models. The DocLayout-YOLO method, which is pre-trained on diverse synthetic document data, significantly outperforms other approaches. This superiority is a key factor in MinerU’s integration of DocLayout-YOLO, contributing to its outstanding overall performance. The table also reveals that, aside from DocLayout-YOLO, other methods perform well on books and academic literature but are less effective on other document types, primarily due to a lack of pre-training on diverse documents.
Table Recognition Results. Table recognition results evaluated by Tree-Edit-Distance-based Similarity (TEDS) metric are presented in Table 7. We evaluate table recognition models across three dimensions on our OmniDocBench table subset: language diversity, table frame types, and special situations. Among all models, OCR-based models demonstrate superior overall performance, with RapidTable achieving the highest scores in language diversity and maintaining stable performance across different frame types. Expert VLMs show competitive results in specific scenarios, with StructEqTable [48] excelling in no frame tables and showing better rotation robustness. General VLMs (Qwen2-VL-7B and InternVL2-8B) exhibit relatively lower but consistent performance, suggesting that while general-purpose VLMs have made progress in table understanding, they still lag behind specialized solutions.
Text Recognition Results. In the traditional OCR task, Table 8 shows that PaddleOCR leads the field, surpassing other models significantly, with GOT also performing relatively well. Selecting these two methods for the OCR module is a prudent choice.
Formula Recognition Results. For formula recognition, the CDM metric provides a clear comparison in Table 9. GPT-4o, Mathpix, and UniMERNet achieve results of 86.8%, 86.6%, and 85.0%, respectively. Notably, GPT-4o excels with a recall rate of 65.5% under strict conditions requiring perfect character accuracy. Although Mathpix shows high character-level precision, it occasionally omits punctuation, such as commas, leading to a lower overall correctness rate. Nonetheless, all three models are strong candidates for formula recognition tasks.
6 Conclusion
This paper addresses the lack of diverse and realistic benchmarks in document parsing research by introducing OmniDocBench, a dataset featuring a variety of page types with comprehensive annotations, along with a flexible and reliable evaluation framework. OmniDocBench enables systematic and fair assessments of document parsing methods, providing crucial insights for advancing the field. Its task-specific and attribute-level evaluations facilitate targeted model optimization, promoting more robust and effective parsing solutions.
Supplementary Material
I More End-to-End Evaluation Results
Table S1 presents the evaluation results of End2End Tables grouped by Table Attributes. As it shows, most of the models perform better in English Tables rather than Chinese ones. Most models perform relatively poorly with Full Frame and No Frame tables. The accuracy of most models is affected by special conditions. Merged cells and formulas mainly test the breadth of data the model can recognize, while colored backgrounds and table rotation test their robustness. The results show that table rotation significantly impacts the accuracy of all models. Pipeline Tools perform well on more challenging tables, but colored backgrounds can affect recognition accuracy. Several Vision Language Models (VLMs) tend to perform worse on tables with merged cells, but colored backgrounds do not significantly impact table recognition accuracy.
Table S2 shows the evaluation results of End2End Text blocks grouped by Text Attributes. Almost all models have lower recognition accuracy in Chinese compared to English. Some models, such as MinerU and Marker, experience a further decrease in accuracy when recognizing mixed Chinese and English content. Complex background colors significantly affect the recognition accuracy of pipeline tools, but they have little impact on VLMs.
Model Type Model Language Table Frame Type Special Situation
EN ZH Mixed Full Omission Three Zero Merge Cell(+/-) Formula(+/-) Colorful (+/-) Rotate(+/-)
Pipeline Tools MinerU 75.7 59.9 79.6 60.0 72.8 70.1 60.4 64.1/66.0 66.7/65.0 59.8/68.1 2.9/66.4
Marker 52.5 43.0 44.2 41.8 55.3 47.1 52.4 43.8/47.0 42.9/46.6 44.3/46.7 6.3/46.6
Mathpix 76.1 64.3 71.9 68.3 79.3 67.0 25.8 71.2/66.4 69.8/67.6 60.5/71.8 20.7/68.8
Expert Vision Models GOT-OCR 51.9 47.0 49.4 46.2 49.3 51.6 47.2 46.5/49.7 46.4/49.1 40.2/52.7 0.0/49.4
Nougat 36.5 0.4 0.0 6.3 3.6 22.2 0.0 15.1/9.1 21.2/8.9 2.8/15.3 0.0/11.4
Vision Language Models GPT4o 71.8 58.8 57.9 63.3 69.5 61.9 31.8 57.5/65.5 61.6/62.9 62.0/63.0 14.5/63.5
Qwen2-VL 57.4 62.9 72.7 70.7 64.1 48.3 57.6 49.4/68.2 48.5/64.7 63.5/60.7 41.6/61.9
InterVL2 61.5 59.3 65.9 59.7 66.5 58.7 56.2 49.6/65.9 54.4/61.6 59.4/60.6 7.3/61.1
Table S1: End-to-End Table TEDS Result grouped by Table Attributes
Model Type Model Language Text background
EN ZH Mixed White Single Multi
Pipeline Tools MinerU 0.123 0.206 0.742 0.163 0.147 0.513
Marker 0.267 0.389 0.499 0.339 0.389 0.497
Mathpix 0.173 0.774 0.538 0.675 0.554 0.570
Expert Vision Models GOT-OCR 0.251 0.763 0.266 0.669 0.595 0.440
Nougat 0.587 0.991 0.983 0.874 0.935 0.972
Vision Language Models GPT4o 0.170 0.647 0.322 0.536 0.423 0.406
Qwen2-VL 0.337 0.575 0.310 0.537 0.400 0.233
InternVL2 0.418 0.606 0.251 0.589 0.366 0.221
Table S2: End-to-End Text Normalized Edit Distance results grouped by Text Attributes. “Mixed” represents a mixture of Chinese and English, “Single” and “Multi” represent single color and multi color.
II Dataset Statistics and Visualization
OmniDocBench contains 981 pages, including 9 types of PDF pages, 4 types of layouts, and 3 types of languages. Some pages also include special conditions, such as watermarks. Table S3 and Figure S1 show the number of pages with each page attribute. Figures S3, S4, S5 and S6 are examples of PDF pages with different PDF types, Layout Types, and Special Issues.
Table S6 and Figure S2 show all annotation categories included in OmniDocBench. All of them are annotated by bounding boxes. There are 15 types of block-level annotations and 4 types of span-level annotations, with span-level annotations nested within the block-level ones. In addition, there are 3 types of annotations marked as page interference information (No.20-22), whose bounding boxes are used to mask the specific regions of the PDF pages to avoid affecting the evaluation results. The recognition annotations are also provided for each annotation category except for Figures. Formulas is written in LaTeX format and Table is annotated in both HTML and LaTeX formats. Others are annotated in plain text.
Furthermore, the Text Attributes are also annotated for each block-level category that contains text. There are 3 types of Text Attributes that might influent OCR accuracy: Language, Text Background Color, and Text Rotation. Table S5 shows the statistics of annotations with specific text attributes. There are 23,010 block-level annotations are labeled with text attributes.
Tables are also annotated with Table Attributes. There are 6 types of Table Attributes that might influent the Table Recognition accuracy: Language, Table Frame Type, Merge Cell, Colorful Background, Contain Formula, and Rotation. Table S5 shows the numbers of annotations with specific table attributes. Figures S7 and S8 are the examples of Tables with different Frames and Special Issues.
III Model Results Visualization
Figures S10, S11, S12, S13, S14, S9, S15, S16 and S17 show the examples of Good model outputs and Bad model outputs of Document Parsing among different PDF types. As it shown, different models exhibit varying performance across different PDF types. For example, MinerU detects all handwritten notes as figures, resulting in very low recognition accuracy in Notes. Marker and InternVL2 experience missed detections, leading to lower scores. InternVL2 and Qwen2-VL, in specific PDF types (such as slides or financial reports), tend to merge multi-column text.
Figures S20, S18 and S19 show the examples of Good model outputs and Bad model outputs under special issues of the PDF pages. It shows that Marker tends to generate typos when the PDF pages are fuzzy scanned or with watermarks, while GOT-OCR fails to recognize content on pages with colored backgrounds. MinerU performs well under special situations, while Mathpix occasionally generates typos.
Figures S21, S22, S23 and S24 show examples of Good model outputs and Bad model outputs for PDF pages with different layouts. MinerU has a low reading order score for single-column layouts primarily because most notes are single-column, and MinerU performs poorly in recognizing Notes, leading to a low reading order score accordingly. InternVL2 scores high in Single-Column layouts but scores poorly on Double-Column and Three-Column layouts. It is mainly due to frequent missed content recognition and errors in reading order judgment in multi-column layouts pages. MinerU’s reading order and recognition accuracy decrease with complex layouts, primarily because it incorrectly merges multiple columns during recognition.
Figures S27 and S28 show the model’s recognition ability under special issues of text. In text recognition with complex background colors, Marker may produce errors or miss content, whereas Qwen2-VL still performs well. Most models fail to recognize text when it is rotated 270 degrees. Some vision language models generate hallucinated information based on the content they can recognize.
Figures S29, S30, S31 and S32 show the examples of good and bad model results for tables with different attributes. For three-line tables, RapidTable demonstrates a good performance with accurate structure recognition, while PaddleOCR shows limitations by missing the last column in its outputs. Interestingly, in tables without frames, PaddleOCR performs well with accurate table predictions, while Qwen2-VL-7B exhibits errors in the last two columns. This indicates that the presence or absence of table frames can significantly impact different models’ performance in different ways. Rotated tables prove to be particularly challenging, with most models, including GOT-OCR, failing to recognize the table structure. However, StructEqTable shows promising results by correctly identifying most of the table content, though with a few detail errors. For tables containing formula, Qwen2-VL-7B shows more accurate table structure recognition compared to InternVL2-8B.
IV Model Settings
For pipeline tools such as MinerU, Marker, and Mathpix, default settings are used for evaluation. Specifically, MinerU with Version 0.9.3555https://github.com/opendatalab/MinerU/releases/tag/magic_pdf-0.9.3-released is employed. For Marker, Version 0.2.17666https://github.com/VikParuchuri/marker/releases/tag/v0.2.17 is evaluated. For Nougat, we utilize its 0.1.0-base model (350M). For GOT-OCR, we employ its format OCR mode to output structured data. For general VLMs, we used the GPT4o, Qwen2-VL-72B, and InternVL2-Llama3-76B by setting the do_sample===False to ensure the reproducibility.
Category Attribute Name Count
PDF Type Book 104
PPT2PDF 133
Research Report 81
Colorful Textbook 96
Exam Paper 114
Magazine 97
Academic Literature 129
Notes 116
Newspaper 111
Layout Type Single Column 477
Double Column 126
Three Column 45
One&More Mixed 120
Complex Layout 213
Language English 290
Simplified Chinese 612
Mixed 79
Special Issues Fuzzy Scan 28
Watermark 65
Colorful Background 246
Table S3: The Page Attributes Statistics of OmniDocBench.
Attribute Category Category Name Count
Language English 5857
Simplified Chinese 16073
EN&CH Mixed 1080
Text Background White 19465
Single-Colored 1116
Multi-Colored 2429
Text Rotate Normal 22865
Rotate90 14
Rotate270 58
Horizontal 421
Table S4: Text Attributes Statistics of OmniDocBench.
Attribute Category Category Name Count
Language English 128
Simplified Chinese 285
EN&CH Mixed 15
Table Frame Type Full Frame 205
Omission Line 62
Three Line 147
No Frame 14
Special Issues Merge Cell 150
Colorful Background 142
Contain Formula 81
Rotate 7
Table S5: Table Attributes Statistics of OmniDocBench.
No. Category Name Explaination Total
1 Title Include main titles, chapter titles, etc. 2972
2 Text Block Text paragraphs, which are usually separated by double line breaks in Markdown. 15979
3 Figure Including images, visual charts, etc. 989
4 Figure Caption Typically starts with ’Figure’ followed by a number, or just descriptive language below the figure. 651
5 Figure Footnotes Descriptive language, apart from the figure caption, usually starts with an asterisk (*). 133
6 Table Content organized in table form usually includes borders or a clear table structure. 428
7 Table Caption Typically starts with ’Table’ followed by a number, or just descriptive language above the Table. 299
8 Table Footnotes Descriptive language, apart from the table caption, usually starts with an asterisk (*). 132
9 Header Information located at the top of a PDF page or in the sidebar, separate from the main content, typically includes chapter names and other details. 1271
10 Footer Information located at the bottom of a PDF page, separate from the main content, typically includes the publisher’s name and other details. 541
11 Page Number It is usually represented by numbers, which may be located at the top, in the sidebar, or at the bottom of the page. 669
12 Page Footnote It provides further explanation of the footnotes marked within the page content. For example, information about the authors’ affiliations. 92
13 Code Block In Markdown, a code block is typically defined using triple backticks (“‘). 13
14 Code Block Caption Descriptive language above the Code Block. /
15 Reference Typically found only in academic literature. 260
16 Text Span Span-Level text box, which is the plain text content can be directly written in Markdown format. 73143
17 Equation Inline Formulas that need to be represented using LaTeX format and embedded within the text. 4009
18 Equation Ignore Some formulas that can be displayed correctly without using LaTeX formatting, such as 15 kg. 3685
19 Footnote Mark Typically embedded within the text as superscripts or subscripts, and their numbering usually corresponds to page footnotes. 357
20 Other Abandoned Categories (Masked) Some uncategorizable, irrelevant page information, such as small icons, etc. 538
21 Masked Text Block (Masked) Some difficult-to-recognize information that disrupts text flow, such as pinyin annotations above Chinese characters. 34
22 Organic Chemical Formula (Masked) Organic chemistry formulas, which are difficult to write using Markdown and are easily recognized as Figures. 24
Table S6: Annotation Explanations and Statistics.
Figure S1: The Data Proportion of Pages for each Attribute in OmniDocBench.
Figure S2: The Visualization of vary Annotations in OmniDocBench.
Figure S3: The Examples of Academic Papers, Books, Textbooks, Notes, and Magazines in OmniDocBench.
Figure S4: The Examples of Finacial Reports, Newspapers, Example Papers, and Slides in OmniDocBench.
Figure S5: The Examples of PDF pages with different Layout Types in OmniDocBench.
Figure S6: The Examples of PDF pages under Special Issues in OmniDocBench.
Figure S7: The Examples of Tables with different Frame in OmniDocBench.
Figure S8: The Examples of Tables under Special Issues in OmniDocBench.
Figure S9: The Good Model Result and Bad Model Result for Academic Papers.
Figure S10: The Good Model Result and Bad Model Result for Books.
Figure S11: The Good Model Result and Bad Model Result for Exam Papers.
Figure S12: The Good Model Result and Bad Model Result for Magazines.
Figure S13: The Good Model Result and Bad Model Result for Newspaper.
Figure S14: The Good Model Result and Bad Model Result for Handwriting Notes.
Figure S15: The Good Model Result and Bad Model Result for Financial Reports.
Figure S16: The Good Model Result and Bad Model Result for Slides.
Figure S17: The Good Model Result and Bad Model Result for Textbooks.
Figure S18: The Good Model Result and Bad Model Result for Fuzzy Scan Pages.
Figure S19: The Good Model Result and Bad Model Result for Pages with Watermark.
Figure S20: The Good Model Result and Bad Model Result for Colorful Background Pages.
Figure S21: The Good Model Result and Bad Model Result for Single Column Pages.
Figure S22: The Good Model Result and Bad Model Result for Double Column Pages.
Figure S23: The Good Model Result and Bad Model Result for Three Column Pages.
Figure S24: The Good Model Result and Bad Model Result for Complex Layout Pages.
Figure S25: The Good Model Result and Bad Model Result for Text Language in Chinese.
Figure S26: The Good Model Result and Bad Model Result for Text Language in English.
Figure S27: The Good Model Result and Bad Model Result for Text with Colorful Background.
Figure S28: The Bad Model Result for Text with Rotation.
Figure S29: The Good Model Result and Bad Model Result for Three Line Frame Table.
Figure S30: The Good Model Result and Bad Model Result for No Frame Table.
Figure S31: The Good Model Result and Bad Model Result for Rotated Table.
Figure S32: The Good Model Result and Bad Model Result for Table with Formula.

### v2

@@
 Abstract
-Document content extraction is crucial in computer vision, especially for meeting the high-quality data needs of large language models (LLMs) and retrieval-augmented generation (RAG) technologies.
-However, current document parsing methods suffer from significant limitations in terms of diversity and comprehensive evaluation.
-To address these challenges, we introduce OmniDocBench, a novel multi-source benchmark designed to advance automated document content extraction.
-OmniDocBench includes a meticulously curated and annotated high-quality evaluation dataset comprising nine diverse document types, such as academic papers, textbooks, slides, among others.
-Our benchmark provides a flexible and comprehensive evaluation framework with 19 layout category labels and 14 attribute labels, enabling multi-level assessments across entire datasets, individual modules, or specific data types.
-Using OmniDocBench, we perform an exhaustive comparative analysis of existing modular pipelines and multimodal end-to-end methods, highlighting their limitations in handling document diversity and ensuring fair evaluation.
-OmniDocBench establishes a robust, diverse, and fair evaluation standard for the document content extraction field, offering crucial insights for future advancements and fostering the development of document parsing technologies.
-The codes and dataset is available in https://github.com/opendatalab/OmniDocBench.
+Document content extraction is a critical task in computer vision, underpinning the data needs of large language models (LLMs) and retrieval-augmented generation (RAG) systems.
+Despite recent progress, current document parsing methods have not been fairly and comprehensively evaluated due to the narrow coverage of document types and the simplified, unrealistic evaluation procedures in existing benchmarks.
+To address these gaps, we introduce OmniDocBench, a novel benchmark featuring high-quality annotations across nine document sources, including academic papers, textbooks, and more challenging cases such as handwritten notes and densely typeset newspapers.
+OmniDocBench supports flexible, multi-level evaluations—ranging from an end-to-end assessment to the task-specific and attribute-based analysis—using 19 layout categories and 15 attribute labels.
+We conduct a thorough evaluation of both pipeline-based methods and end-to-end vision-language models, revealing their strengths and weaknesses across different document types.
+OmniDocBench sets a new standard for the fair, diverse, and fine-grained evaluation in document parsing.
+Dataset and code are available at https://github.com/opendatalab/OmniDocBench.
 footnotetext: The authors contributed equally.footnotetext: Project lead.footnotetext: Corresponding author (heconghui@pjlab.org.cn).
 1 Introduction
-Figure 1: OmniDocBench Data Diversity.
-It contains 9 PDF page types, along with Layout Annotations and Recognition Annotations.
-Furthermore, there are 5 Page Attributes, 3 Text Attributes, and 6 Table Attributes.
-Document parsing is a foundational task in computer vision, focused on accurately extracting content from documents [18, 36, 39, 41, 45].
-High-quality document content extraction typically involves the integration of multiple algorithmic modules.
-Layout detection algorithms identify different content areas on a page, OCR technology converts images of text regions into text, while formula and table recognition models identify specific regions and transform them into corresponding source code.
-These modules and reading order algorithms form a comprehensive process of converting documents into machine-readable formats.
-With large models increasingly requiring high-quality data, the importance of document content extraction has become more pronounced.
-Although vast amounts of data are available online for training, knowledge-rich document data is relatively scarce.
-Documents such as academic papers and technical reports contain rich structured information that can significantly enhance the knowledge depth of large models.
-Moreover, the development of retrieval-augmented generation (RAG) [21, 10] technology relies on extracting accurate information from documents to improve the quality and relevance of generated content.
-Consequently, research in document content extraction has intensified, leading to a series of pipeline-based high-quality document extraction algorithms [36] and the emergence of end-to-end multimodal large model solutions [5, 40, 27, 39, 3, 6, 42].
-These methods have significantly improved document content parsing quality, providing robust support for the needs of large models and RAG technology.
-In analyzing current module-based pipeline and multimodal end-to-end methods, we identified several limitations.
-For instance, methods like Marker and MinerU, which are mainstream pipeline methods, primarily evaluate individual modules on academic paper data, lacking document diversity and comprehensive evaluation results.
-Although MinerU considers the generalization of diverse data, it only demonstrates this through a single model and visualization results, lacking overall end-to-end evaluation.
-Multimodal large model methods [5, 40, 27, 39, 3], while easier to use than pipeline methods, lack performance validation on diverse documents, and some evaluation metrics are inadequate.
-Additionally, these methods often use data similar to their training set distribution for comparison, resulting in unfair evaluations.
-Overall, current document content extraction faces the following challenges:
-  • Limited document types.
-Current evaluations mostly focus on a single type of academic paper, while real-world scenarios include textbooks, exam papers, financial reports, newspapers, magazines, and other document types.
-  • Monotonous evaluation dimensions.
-Pipeline-based methods typically evaluate specific algorithmic modules, such as OCR, layout detection, or formula recognition, while the overall quality of parsing results requires comprehensive metrics.
-  • Inadequate evaluation metrics.
-Multimodal large model approaches attempt to evaluate document parsing quality across multiple dimensions, such as dividing document content into text, formulas, tables, etc.
-However, these models commonly employ evaluation metrics such as BLEU scores or Edit distances, which fail to accurately and fairly assess parsing effectiveness when dealing with markup languages like LaTeX or HTML that allow diverse syntactic expressions.
-Building a diverse, comprehensive, and accurate evaluation system poses significant challenges, requiring diverse and high-quality data annotation and reasonable evaluation metrics.
-While READOC extends the evaluation scope to include GitHub README files based on Nougat, there remains a substantial gap in real-world diversity, and the evaluation dimensions lack consideration of attributes.
-In contrast, this paper proposes a document content extraction benchmark, OmniDocBench, characterized by diverse types, detailed annotations, and reasonable evaluation (Figure 1).
-The specific contributions are as follows:
-  • High-quality, diverse evaluation set: Through automated annotation, manual verification, and expert review, we construct a comprehensive, detailed, high-quality OmniDocBench evaluation set, encompassing nine types of diverse document pages, including papers, textbooks, exam questions, and research reports.
-  • Flexible and comprehensive evaluation dimension support: The OmniDocBench validation set covers 19 layout category labels and 14 attribute labels.
-To facilitate user evaluation from an overall, single module, or different data types, we provide end-to-end evaluation, single algorithm module evaluation, and attribute-based evaluation, covering various evaluation needs.
-  • Comprehensive evaluation of mainstream methods: Based on OmniDocBench, we conduct a comprehensive evaluation of current mainstream modular pipeline and end-to-end large model methods, providing a fairer assessment of existing methods and summarizing the shortcomings of current document parsing methods, thereby guiding further development in document parsing.
+Figure 1: Results of End-to-End Text Recognition on OmniDocBench across 9 PDF page types.
+Figure 2: Overview of OmniDocBench Data Diversity.
+The benchmark includes 9 diverse PDF document types.
+It supports rich annotation types, including layout annotations (e.g., title, table, figure) and recognition annotations (e.g., text spans, equations, tables).
+Each page is annotated with 6 page-level attributes (e.g., PDF type, layout type), along with fine-grained 3 text attributes (e.g., language) and 6 tables attributes (Items under “special issues” are treated as individual binary attributes (yes/no)), enabling detailed and robust evaluation.
+As large language models [1, 39, 44, 28] increasingly rely on high-quality, knowledge-rich data, the importance of accurate document parsing has grown substantially.
+Document parsing, a core task in computer vision and document intelligence, aims to extract structured, machine-readable content from unstructured documents such as PDFs.
+This task is particularly critical for ingesting academic papers, technical reports, textbooks, and other rich textual sources into large language models, thereby enhancing their factual accuracy and knowledge grounding [19, 42, 45, 47, 52].
+Moreover, with the emergence of retrieval-augmented generation (RAG) systems [22, 12], which retrieve and generate answers conditionally with external documents, the demand for precise document understanding has further intensified.
+To address this challenging task, two main paradigms have emerged: 1) Pipeline-based approaches that decompose the task into layout analysis, OCR, formula/table recognition, and reading order estimation [42, 34]; and 2) End-to-end vision-language models (VLMs) that directly output structured representations (e.g., Markdown) [7, 46, 29, 45, 3, 8, 48].
+Although both approaches have demonstrated promising results, conducting a broad comparison of their effectiveness remains challenging due to the absence of a comprehensive and unified evaluation benchmark.
+As shown in Table 1, for pipeline-based document parsing systems, dedicated benchmarks [26, 54, 10] have been developed to target specific sub-tasks.
+For end-to-end evaluation, works like Nougat [7] and GOT-OCR [45] provide relatively small validation sets and assess predictions using page-level metrics such as Edit Distance [21].
+However, these benchmarks present several key limitations: 1) Limited document diversity: Existing datasets primarily focus on academic papers, overlooking other real-world document types such as textbooks, exams, financial reports, and newspapers; 2) Inconsistent evaluation metrics: Current benchmarks rely heavily on generic text similarity metrics (e.g., Edit Distance [21] and BLEU [33]), which fail to fairly assess the accuracy of formulas and tables in LaTeX or HTML formats that allow for diverse syntactic expressions; and 3) Lack of fine-grained evaluation: Most evaluations report only an overall score, lacking insights into specific weaknesses, such as element-level score (e.g., text vs.
+formula) or per document-type performance (e.g., magazine or notes).
+To address these limitations, we introduce OmniDocBench, a new benchmark designed to provide a rigorous and comprehensive evaluation for document parsing models across both pipeline-based and end-to-end paradigms.
+In summary, our benchmark introduces the following key contributions:
+  • High-quality, diverse evaluation set: We include pages from 9 distinct document types, ranging from textbooks to newspapers, annotated using a combination of automated tools, manual verification, and expert review.
+  • Flexible, multi-dimensional evaluation: We support comprehensive evaluation at three levels—end-to-end, task-specific, and attribute-based.
+End-to-end evaluation measures the overall quality of full-page parsing results.
+Task-specific evaluation allows users to assess individual components such as layout detection, OCR, table recognition, or formula parsing.
+Attribute-based evaluation provides fine-grained analysis across 9 document types, 6 page-level attributes and 9 bbox-level attributes.
+  • Comprehensive benchmarking of state-of-the-art methods: We systematically evaluate a suite of representative document parsing systems, including both pipeline-based tools and VLMs, providing the most comprehensive comparison and identifying performance bottlenecks across document types and content structures.
 2 Related Work
-Benchmark Document Categories Annotaion Type Single-Module Eval End-to-End Eval
-BBox Text Table Order Formula OCR DLA TR MFR OCR ROD TR MFR
-Single-Module Eval Benchmark
+Benchmark Document Domain Annotaion Type Single-Task Eval End-to-End Eval
+BBox Text Table Formula Attributes OCR DLA TR MFR OCR TR MFR ROD
+Single-Task Eval Benchmark
 Robust Reading [20] 1
 PubLayNet [49], DocBank [26], DocLayNet [35], M6Doc [9] 1, 1, 5, 6
-PubTabNet [47],TableX [9], TableBank [23] 1, 1, 1
+PubTabNet [54],TableX [11] 1, 1
+TableBank [25] 1
 Im2Latex-100K [10],UniMER-Test [40] 1
 End-to-end Eval Benchmarks
+Fox [29] 2
 Nougat [7] 1
-Fox [27] 2
 GOT OCR 2.0 [45] 2
-READoc [26] 2
 OmniDocBench 9
-Table 1: A comparison between OmniDocBench and existing DCE benchmarks.
-OCR: Optical Character Recognition; DLA: Document Layout Analysis; MFR: Math Formula Recognition; TR: Table Recognition; ROD: Reading Order Detection
-2.1 Traditional Document Content Extraction
-Document Content extraction remains a challenging task, and there is yet to emerge a unified benchmark tailored for real-world scenarios.
-Traditional algorithms typically employ multiple expert modules to handle different extraction subtasks, such as document layout detection [16, 46, 32, 11], optical character recognition (OCR) [33, 22, 28, 14, 37], formula recognition [44, 25, 4, 34], and table recognition [15, 17, 22].
-While expert models of these subtasks are advancing rapidly, recent work such as Mineru [36] attempts to concatenate multiple expert modules into a pipeline and provides a high-precision open-source solution for document content extraction.
-READOC [26] also unifies heterogeneous evaluation methods from the perspective of Document Structure Extraction, breaking down texts, images, formulas, tables, and other dimensions for evaluation, thus offering a solution-oriented towards real-world scenarios for DSE tasks.
-However, due to the complexity of Document data sources and the intricacies of PDF document information, previous efforts still fall short in terms of data diversity, failing to cover the categories users encounter in practical applications.
-Similarly, there is an issue with the explainability of document parsing.
+Table 1: A Comparison between OmniDocBench and existing benchmarks.
+BBox: Bounding boxes.
+Text: Text in Unicode.
+Table: Table in LaTeX/HTML/Markdown.
+Formula: Formula in LaTeX.
+Attributes: Page- and BBox-Level Attributes.
+OCR: Optical Character Recognition; DLA: Document Layout Analysis; TR: Table Recognition; MFR: Math Formula Recognition; ROD: Reading Order Detection
+2.1 Pipeline-based Document Content Extraction
+Pipeline-based methods treat the document content extraction task as a collection of single modules, such as document layout detection [17, 53, 36, 13], optical character recognition [38, 23, 30, 15, 43], formula recognition [51, 27, 6, 40], and table recognition [16, 18, 23].
+In this sense, such methods can utilize different expert models to address each specific task.
+Marker [34] integrates open-source models to parse documents into structured formats such as Markdown, JSON, and HTML.
+To get higher accuracy, an optional LLM-enabled version can also be integrated to merge tables across pages, handle inline math, and so on.
+Similarly, MinerU [42] first utilizes a layout detection model to segment the document page into different regions, then applies task-specific models for corresponding regions.
+Finally, it outputs the complete content in Markdown format with a reading order algorithm.
+By leveraging lightweight models and parallelized operations, pipeline-based methods can achieve efficient parsing speeds.
 2.2 VLM-based Document Content Extraction
-The emergence of Vision-Language Models (VLMs) [38, 6, 1, 12] has revolutionized the field of document content extraction.
-These models leverage multi-modality capability to achieve remarkable performance in document understanding tasks.
-Document extraction tools powered by VLMs excel at comprehending both visual layouts and textual content, effectively handling complex document structures while capturing rich contextual information.
-Representative works such as Nougat [5], Vary [40], Fox [27], and GOT [39], along with recent advances [13, 29], demonstrate significant progress in automated document parsing and comprehension.
-Despite these advances, the field lacks a standardized and unified benchmark for evaluating VLM-based document extraction task.
-This absence has hindered objective assessment of PDF document processing capabilities and impeded fair comparison across different approaches.
-To address this limitation, we present OmniDocBench, a comprehensive end-to-end benchmark designed specifically for evaluating VLM-based document parsing in real-world scenarios.
-2.3 Benchmark for Document Content Extraction
-An end-to-end benchmark for PDFs can intuitively reflect the effectiveness of PDF extraction tools, which is crucial for their iteration and selection.
-However, current benchmarks predominantly focus on module-level evaluations; we have listed related benchmarks in Table 1.
-Additionally, while there are existing end-to-end benchmarks, they lack detailed annotation rules and suffer from insufficient diversity, as well as unreasonable metrics for formula and table evaluations.
-For example, READOC [26] covers only two types of sources—arXiv and GitHub—and uses EDS [20] and TEDS [47] to compute metrics for formulas and tables, which may lead to inaccuracies CDM [35].
-Therefore, there is a need for a more finely annotated, diverse, and reasonably evaluated end-to-end benchmark.
+Document understanding and optical character recognition (OCR) are crucial tasks for evaluating the perception capabilities of vision-language models (VLMs).
+By incorporating extensive OCR corpus into the pretraining stage, VLMs like GPT4o [2] and Qwen2-VL [3] have demonstrated comparable performance in document content extraction tasks.
+Unlike pipeline-based methods, VLMs perform document parsing in an end-to-end manner.
+Furthermore, without requiring specialized data fine-tuning, these models are able to deal with diverse and even unseen document types for their generalization capabilities.
+To integrate the efficiency of lightweight models and the generalizability of VLMs, many works [7, 46, 29, 45, 14, 32] have focus on training specialized end-to-end expert models for document parsing.
+These VLM-driven models excel at comprehending both visual layouts and textual contents, balancing a trade-off between accuracy and efficiency.
+2.3 Benchmarks for Document Content Extraction
+Document content extraction requires the ability to understand document layouts and recognize various types of content.
+However, current benchmarks fall short of a comprehensive page-level evaluation, as they focus solely on evaluating the model’s performance on module-level recognition.
+PubLayNet [49] and concurrent benchmarks [26, 35, 9] specialize in evaluating a model’s ability to detect document page layouts.
+OCRBench [31] proposes five OCR-related tasks with a greater emphasis on evaluating the model’s visual understanding and reasoning capabilities.
+Only line-level assessments are provided for text recognition and handwritten mathematical expression recognition (HMER).
+Similarly, single-module benchmarks [20, 54, 26, 40] disentangle the task into different dimensions and focus narrowly on specific parts.
+Such paradigm overlooks the importance of structural and semantic information like the reading order and fails to evaluate the model’s overall ability when processing the full-page documents as a whole.
+Page-level benchmarks have been proposed alongside some recent VLM-driven expert models [29, 7, 45].
+However, the robustness of these benchmarks is compromised by limitations in data size, language, document type, and annotation.
+For example, Nougat [7] evaluates models using only printed English documents collected from arXiv while the page-level benchmark introduced by GOT-OCR [45] consists of only 90 pages of Chinese and English documents in total.
+Commonly-seen document types like handwritten notes, newspapers, and exam papers are further neglected.
+Lacking detailed annotations, the benchmarks can only conduct naive evaluation between the full-page results of Ground Truths and predictions without special handling for different output formats and specialized metrics for different content types.
+The evaluation of the model performance can be severely biased due to limited document domains, unaligned output format and mismatched metrics.
+Therefore, there is an urgent need for a more finely annotated, diverse, and reasonable page-level document content extraction benchmark.
 3 OmniDocBench Dataset
-Constructing a diverse and comprehensive document parsing benchmark with precise annotations is a formidable challenge.
+Constructing a diverse and comprehensive document parsing benchmark with precise annotations is a significant challenge.
 As illustrated in Figure 3, we have designed a systematic and professional annotation framework for OmniDocBench, encompassing data acquisition, intelligent pre-annotation, and manual refinement.
@@
   • Page Diversity. We sourced document pages from a variety of origins to ensure a wide range of document types.
-  • Comprehensive Annotation. We meticulously annotated all elements on the pages, including bounding boxes, specific content, and various potential attributes.
+  • Comprehensive Annotation. We meticulously annotated all elements on the pages, including bounding boxes, specific contents, and various potential attributes.
   • Annotation Accuracy. By integrating semi-automated annotation processes, annotator corrections, and expert quality checks, we ensure the reliability of all annotations.
@@
 During the data acquisition phase, we sourced document pages from diverse origins and used clustering algorithms to initially select visually diverse pages, followed by manual annotation of page attributes to finalize the OmniDocBench pages.
-Specifically, we collected 200,000 initial PDF documents from Common Crawl, Google, Baidu search engines, and internal data.
+Specifically, we collected over 200,000 initial PDF documents from Common Crawl, Google, Baidu search engines, and internal data.
 Subsequently, we extracted visual features from these document pages using ResNet-50 and performed clustering using Faiss  [^1] https://github.com/facebookresearch/faiss, sampling 6,000 visually diverse pages from 10 cluster centers.
 Finally, annotators provided page-level attribute annotations, including page type, layout type, and language type, and further balanced the selection to 981 samples for the final dataset.
-The OmniDocBench dataset includes pages with 9 types of pages, multiple layout categories, and various attribute annotations, covering a wide range of real-world scenarios.
+The OmniDocBench dataset includes pages from nine distinct types, multiple layout categories, and various attribute annotations, covering a wide range of real-world scenarios.
 Figure 3: Overview of the OmniDocBench dataset construction.
@@
 3.2.1 Annotation Types
-Layout Detection Annotations: Unlike typical layout detection tasks, OmniDocBench includes four comprehensive types of annotations: (1) Layout Bounding Box Annotations: Locating information for 19 types of regions such as titles, text paragraphs, tables, and images.
-(2) Layout Attribute Annotations: Detailed attribute annotations for detected boxes, including 3 text box attributes, 6 table attributes, and 2 formula attributes.
+Layout Detection Annotations: Unlike typical layout detection tasks, OmniDocBench includes four comprehensive types of annotations: (1) Layout Bounding Box Annotations: Positioanl information for 19 distinct region categories such as titles, text paragraphs, tables, and images.
+(2) Layout Attribute Annotations: Detailed attribute annotations for detected boxes, including 3 text box attribute categories, 6 table attribute categories, 9 bbox-level attribute labels in total.
 (3) Reading Order Annotations: Annotating the reading sequence of detected boxes.
@@
 Similarly, for cross-page paragraphs, we annotate affiliation relationships.
-Content Recognition Annotations: Based on the format of the content area, we conduct the following three types of area annotations: (1) Text Annotations: Pure text annotations for titles, text paragraphs, and other plain text content.
+Content Recognition Annotations: Based on the content type within each region, we conduct the following three types of annotations: (1) Text Annotations: Pure text annotations for titles, text paragraphs, and other plain text content.
 (2) Formula Annotations: LaTeX format annotations for inline formulas, display formulas, and subscripts.
@@
 3.2.2 Annotation Process
-For these annotation tasks on diverse pages, we design a standardized process to ensure quality and efficiency, comprising intelligent pre-annotation, annotator correction, and expert quality inspection.
-Intelligent Pre-Annotation.
+For these annotation tasks on diverse pages, we design a standardized process to ensure quality and efficiency, comprising intelligent automatic annotation, annotator correction, and expert quality inspection.
+Automatic Annotation.
 Manually annotating entire documents is time-consuming and costly.
@@
 Annotator Correction.
-After layout detection phase, annotators refine the detection boxes and enhance annotations with reading order and affiliation details.
+After the layout detection phase, annotators refine the detection boxes and enhance annotations with reading order and affiliation details.
 Each character is verified to ensure accuracy in content recognition.
@@
 Despite thorough annotator corrections, the complexity of formulas and tables may result in residual issues.
-To address these, we use CDM’s rendering techniques to identify unrenderable elements.
-These are then reviewed and corrected by three researchers to ensure accuracy and fidelity in the final annotations.
+To address these, we use CDM’s rendering techniques [41] to identify unrenderable elements.
+These elements are then reviewed and corrected by three researchers to ensure accuracy in the final annotations.
 3.3 Dataset Statistics
@@
 Each page is annotated with global attributes, including text language, column layout type, and indicators for blurred scans, watermarks, and colored backgrounds.
-Annotation Diversity: OmniDocBench contains over 10,0000 annotations for page detection and recognition: (1) More than 20,000 block-level annotations across 15 categories, including over 9,000 text paragraphs, 989 image boxes, 428 table boxes, and so on.
+Annotation Diversity: OmniDocBench contains over 100,000 annotations for page detection and recognition: (1) More than 20,000 block-level annotations across 15 categories, including over 15,979 text paragraphs, 989 image boxes, 428 table boxes, and so on.
 All document components except headers, footers, and page notes are labeled with reading order information, totaling over 16,000 annotations.
-(2) The dataset also includes more than 80,000 span-level annotations across four categories, with 4,000 inter-line formulas and footnote markers represented in LaTeX format, while the remaining annotations are in text format.
+(2) The dataset also includes more than 70,000 span-level annotations across 4 categories, with 4,009 inline formulas and 357 footnote markers represented in LaTeX format, while the remaining annotations are in text format.
 Annotation Attribute Diversity: (1) Text Attributes: All block-level annotations, except for tables and images, include text attribute tags.
-In addition to standard Chinese and English text, there are over 2,000 blocks with complex backgrounds and 146 with rotated text.
-(2) Table Attributes: Besides standard Chinese and English tables, there are 142 with complex backgrounds, 81 containing formulas, 150 with merged cells, and 7 vertical tables.
+In addition to standard Chinese and English text, there are over 2,000 blocks with complex backgrounds and 493 with rotated text.
+(2) Table Attributes: In addition to standard Chinese and English tables, there are 142 tables with complex backgrounds, 81 containing formulas, 150 with merged cells, and 7 vertical tables.
 4 OmniDocBench Evaluation Methodology
 To provide a fair and comprehensive evaluation for various models, we proposed an end-to-end evaluation pipeline consisting of several modules, including extraction, matching algorithm, and metric calculation, as shown in Figure 4.
-It ensures that OmniDocBench automatically performs unified evaluation on end-to-end DCE tasks, thereby producing reliable and effective evaluation results.
+It ensures that OmniDocBench automatically performs unified evaluation on document parsing, thereby producing reliable and effective evaluation results.
 Figure 4: OmniDocBench Evaluation Pipeline.
 4.1 Extraction
-Preprocessing: The model-generated markdown text should be preprocessed, which includes removing images, eliminating markdown tags at the beginning of the document, and standardizing the number of repeated characters.
-Special Component Extraction: Extraction is primarily carried out using regular expression matching.
-To ensure that the extraction of content does not interfere with each other, it is necessary to follow a specific order.
+Preprocessing.
+The model-generated markdown text should be preprocessed, which includes removing images, eliminating markdown tags at the beginning of the document, and standardizing the number of repeated characters.
+Elements Extraction.
+Extraction is primarily carried out using regular expression matching.
+To ensure that the extraction of elements does not interfere with each other, it is necessary to follow a specific order.
 The extraction sequence is as follows: LaTeX tables, HTML tables, display formulas, markdown tables (which are then converted into HTML format), and code blocks.
-Pure Text Extraction: After extracting special components, the remaining content is considered pure text.
+Pure Text Extraction.
+After extracting special components, the remaining content is considered pure text.
 Paragraphs are separated by double line breaks, allowing them to participate in subsequent matching processes, thus aligning with reading order annotation units in the GTs.
@@
 Additionally, previously extracted code blocks are merged into the text category for processing.
-Inline Formula Format Converting: We standardized inline formulas within paragraphs to Unicode format.
+Inline Formula Format Converting.
+We standardized inline formulas within paragraphs to Unicode format.
 This was necessary because different models produce inconsistent outputs for inline formulas.
@@
 Instead, we include them in their Unicode format alongside the text paragraphs for evaluation.
-Reading Order Extraction: Upon completion of the extraction, the start and end positions of the extracted content in the original markdown are recorded for subsequent reading order calculation.
+Reading Order Extraction. Upon completion of the extraction, the start and end positions of the extracted content in the original markdown are recorded for subsequent reading order calculation.
 […]
 EN ZH EN ZH EN ZH EN ZH EN ZH EN ZH EN ZH
-Pipeline Tools MinerU 0.058 0.211 0.278 0.577 66.9 49.5 79.4 62.7 0.305 0.461 0.079 0.288 0.180 0.384
-Marker 0.141 0.303 0.667 0.868 18.4 12.7 54.0 45.8 0.718 0.763 0.138 0.306 0.416 0.560
-Mathpix 0.101 0.358 0.306 0.454 71.4 72.7 77.9 68.2 0.322 0.416 0.105 0.275 0.209 0.376
-Expert VLMs GOT-OCR 0.187 0.315 0.360 0.528 81.8 51.4 53.5 48.0 0.521 0.594 0.141 0.28 0.302 0.429
-Nougat 0.365 0.998 0.488 0.941 17.4 16.9 40.3 0.0 0.622 1.000 0.382 0.954 0.464 0.973
-General VLMs GPT4o 0.144 0.409 0.425 0.606 76.4 48.2 72.8 63.7 0.363 0.474 0.128 0.251 0.265 0.435
-Qwen2-VL 0.252 0.251 0.468 0.572 54.9 60.9 59.9 66.8 0.591 0.587 0.255 0.223 0.392 0.408
-InternVL2 0.353 0.290 0.543 0.701 69.8 49.6 63.8 61.1 0.616 0.638 0.317 0.228 0.457 0.464
+Pipeline Tools MinerU [42] 0.061 0.215 0.278 0.577 57.3 42.9 78.6 62.1 0.18 0.344 0.079 0.292 0.15 0.357
+Marker [34] 0.08 0.315 0.53 0.883 17.6 11.7 67.6 49.2 0.619 0.685 0.114 0.34 0.336 0.556
+Mathpix  [^4] https://mathpix.com/ 0.105 0.384 0.306 0.454 62.7 62.1 77.0 67.1 0.243 0.32 0.108 0.304 0.191 0.365
+Expert VLMs GOT-OCR [45] 0.189 0.315 0.360 0.528 74.3 45.3 53.2 47.2 0.459 0.52 0.141 0.28 0.287 0.411
+Nougat [7] 0.365 0.998 0.488 0.941 15.1 16.8 39.9 0.0 0.572 1.000 0.382 0.954 0.452 0.973
+General VLMs GPT4o [2] 0.144 0.409 0.425 0.606 72.8 42.8 72.0 62.9 0.234 0.329 0.128 0.251 0.233 0.399
+Qwen2-VL-72B [44] 0.096 0.218 0.404 0.487 82.2 61.2 76.8 76.4 0.387 0.408 0.119 0.193 0.252 0.327
+InternVL2-76B [8] 0.353 0.290 0.543 0.701 67.4 44.1 63.0 60.2 0.547 0.555 0.317 0.228 0.44 0.443
 Table 2: Comprehensive evaluation of document parsing algorithms on OmniDocBench: performance metrics for text, formula, table, and reading order extraction, with overall scores derived from ground truth comparisons.
-Model Type Models Book Slides Financial Report Textbook Exam Paper Magazine Academic Papers Notes Newspaper Average
-Pipeline Tools MinerU 0.044 0.124 0.033 0.102 0.159 0.072 0.025 0.984 0.148 0.188
-Marker 0.188 0.327 0.087 0.292 0.423 0.134 0.102 0.470 0.270 0.255
-Mathpix 0.131 0.168 0.202 0.199 0.278 0.138 0.091 0.631 0.648 0.276
-Expert VLMs GOT-OCR 0.105 0.222 0.067 0.132 0.204 0.198 0.179 0.388 0.771 0.252
-Nougat 0.734 0.958 1.000 0.820 0.930 0.83 0.214 0.991 0.871 0.816
-General VLMs GPT4o 0.157 0.163 0.348 0.187 0.281 0.173 0.146 0.607 0.751 0.313
-Qwen2-VL 0.094 0.08 0.145 0.148 0.219 0.065 0.315 0.298 0.79 0.239
-InternVL2 0.216 0.098 0.162 0.184 0.247 0.150 0.419 0.226 0.903 0.289
+Model Type Models Book Slides Financial Report Textbook Exam Paper Magazine Academic Papers Notes Newspaper Overall
+Pipeline Tools MinerU [42] 0.055 0.124 0.033 0.102 0.159 0.072 0.025 0.984 0.171 0.206
+Marker [34] 0.074 0.34 0.089 0.319 0.452 0.153 0.059 0.651 0.192 0.274
+Mathpix 4 0.131 0.22 0.202 0.216 0.278 0.147 0.091 0.634 0.69 0.3
+Expert VLMs GOT-OCR [45] 0.111 0.222 0.067 0.132 0.204 0.198 0.179 0.388 0.771 0.267
+Nougat [7] 0.734 0.958 1.000 0.820 0.930 0.83 0.214 0.991 0.871 0.806
+General VLMs GPT4o [2] 0.157 0.163 0.348 0.187 0.281 0.173 0.146 0.607 0.751 0.316
+Qwen2-VL-72B [44] 0.096 0.061 0.047 0.149 0.195 0.071 0.085 0.168 0.676 0.179
+InternVL2-76B [8] 0.216 0.098 0.162 0.184 0.247 0.150 0.419 0.226 0.903 0.3
 Table 3: End-to-end text recognition performance on OmniDocBench: evaluation using edit distance across 9 PDF page types.
-Models Fuzzy Water Color Mean Variance
-Pipeline Tools
-MinerU 0.15 0.151 0.107 0.136 0.0004
-Marker 0.286 0.436 0.290 0.337 0.0049
-Mathpix 0.294 0.290 0.182 0.255 0.0027
-Expert VLMs
-GOT-OCR 0.175 0.190 0.186 0.184 0.0000
-Nougat 0.934 0.915 0.873 0.907 0.0006
-General VLMs
-GPT4o 0.263 0.195 0.184 0.214 0.0012
-Qwen2-VL 0.101 0.157 0.114 0.124 0.0006
-InternVL2 0.120 0.197 0.155 0.157 0.0010
+Models Fuzzy Water Color None
+MinerU [42] 0.15/0.048 0.151/0.031 0.107/0.052 0.079/0.035
+Marker [34] 0.333/0.092 0.484/0.126 0.319/0.127 0.062/0.125
+Mathpix 4 0.294/0.064 0.290/0.059 0.216/0.09 0.135/0.043
+GOT-OCR [45] 0.175/0.05 0.190/0.056 0.186/0.097 0.177/0.081
+Nougat [7] 0.934/0.051 0.915/0.071 0.873/0.096 0.615/0.208
+GPT4o [2] 0.263/0.078 0.195/0.057 0.184/0.078 0.186/0.072
+Qwen2-VL-72B [44] 0.082 /0.01 0.172/ 0.078 0.104/0.05 0.084/0.042
+InternVL2-76B [8] 0.120/0.013 0.197/0.042 0.155/0.059 0.261/0.082
 Table 4: End-to-end text recognition on OmniDocBench: evaluation under various page attributes using the edit distance metric.
+The value is Mean/Variance of scores in the attribute group.
 Columns represent: Fuzzy (Fuzzy scan), Water (Watermark), Color (Colorful background).
-Models Single Double Three Complex Mean Variance
-Pipeline Tools
-MinerU 0.311 0.101 0.117 0.376 0.226 0.0143
-Marker 0.231 0.251 0.309 0.378 0.292 0.0033
-Mathpix 0.189 0.175 0.225 0.413 0.250 0.0091
-Expert VLMs
-GOT-OCR 0.163 0.145 0.257 0.468 0.258 0.0165
-Nougat 0.852 0.601 0.662 0.873 0.747 0.0139
-General VLMs
-GPT4o 0.109 0.204 0.254 0.426 0.248 0.0132
-Qwen2-VL 0.098 0.248 0.517 0.429 0.323 0.0263
-InternVL2 0.082 0.312 0.682 0.444 0.380 0.0472
-Table 5: End-to-end reading order evaluation on OmniDocBench: results across different column layout types using Normalized Edit Distance.
+None (No special issue)
+Models Single Double Three Complex
+MinerU [42] 0.311/0.187 0.101/0.013 0.117/0.046 0.385/0.057
+Marker [34] 0.299/0.143 0.299/0.299 0.149/0.063 0.363/0.086
+Mathpix 4 0.207/0.123 0.188/0.07 0.225/0.029 0.452/0.177
+GOT-OCR [45] 0.163/0.106 0.145/0.059 0.257/0.072 0.468/0.185
+Nougat [7] 0.852/0.084 0.601/0.224 0.662/0.093 0.873/0.09
+GPT4o [2] 0.109/0.112 0.204/0.076 0.254/0.046 0.426/0.188
+Qwen2-VL-72B [44] 0.066/0.048 0.145/0.049 0.204/0.055 0.394/0.203
+InternVL2-76B [8] 0.082/0.052 0.312/0.069 0.682/0.098 0.444/0.174
+Table 5: End-to-end reading order evaluation on OmniDocBench: results across different column layout types using Normalized Edit Distance. The value is Mean/Variance of scores in the attribute group.
 4.2 Matching Algorithm
-To avoid the impact of paragraph splitting on the final results, we proposed a method, Adjacency Search Match, that merges and splits paragraphs in both GTs and Preds to achieve the best possible match.
+Adjacency Search Match.
+To avoid the impact of paragraph splitting on the final results, we proposed Adjacency Search Match, that merges and splits paragraphs in both GTs and Preds to achieve the best possible match.
 The specific strategy involves: i) Calculate a metrix of Normalized Edit Distance between GTs and Preds.
-If the similarity between a Pred and a GT exceeds a specific threshold, they are considered a successful match.
+The Pred and GT pairs whose similarity exceeds a specific threshold are considered as successful match.
 ii) For the rest, we apply fuzzy matching to determine whether one string is a subset of another string.
-If so, we further apply the truncation and merging algorithm which would try to merge adjacent paragraph.
+If so, we further apply the merging algorithm which would try to merge adjacent paragraph.
 This process would continue to merge more paragraph until the Normalized Edit Distance starts to decrease.
 After this process, the best match will be found for GTs and Preds.
-4.3 Metric Calculation
-Ignore Handling: We implement an ignore logic for certain components in PDF page content, meaning they participate in matching but are excluded from metric calculations.
+Ignore Handling.
+We implement an ignore logic for certain components in PDF page content, meaning they participate in matching but are excluded from metric calculations.
 This is mainly because of inconsistent output standards among models, which should not affect the validation results.
 For fairness, we ignore: (1) Headers, footers, page numbers, and page footnotes, which are handled inconsistently by different models.
-(2) Captions for figures, tables, and footnotes often have uncertain placement, complicating reading order.
+(2) Captions for figures, tables, and footnotes often have uncertain placements, thus complicating the reading order.
 Additionally, some models embed table captions in HTML or LaTeX tables, while others treat them as plain text.
-Metric: Different calculation methods are used for various document components: (1) Pure Text: We calculate Normalized Edit Distance, averaging these metrics at the sample level to obtain the final scores.
-(2) Tables: All tables are converted to HTML format before calculating the TEDS metric and Normalized Edit Distance.
-(3) Formulas: Formulas are currently evaluated using the CDM [35], Normalized Edit Distance, and BLEU.
-We did not convert interline formulas into Unicode because Unicode cannot represent certain complex formulas, such as matrices.
-(4) Reading Order: Reading order use the Normalized Edit Distance as metric.
-It only involves text components, where tables, images, and ignored components do not participate in the final reading order calculation.
-5 Benchmarks
-Model Book Slides Research Report Textbook Exam Paper Magazine Academic Literature Notes Newspaper Average mAP
-DiT-L 43.44 13.72 45.85 15.45 3.40 29.23 66.13 0.21 23.65 26.90
-LayoutLMv3 42.12 13.63 43.22 21.00 5.48 31.81 64.66 0.80 30.84 28.84
-DOCX-Chain 30.86 11.71 39.62 19.23 10.67 23.00 41.60 1.80 16.96 21.27
-DocLayout-YOLO 43.71 48.71 72.83 42.67 35.40 51.44 66.84 9.54 57.54 48.71
+Model Backbone Params Book Slides Research Report Textbook Exam Paper Magazine Academic Literature Notes Newspaper Average
+DiT-L [24] ViT-L 361.6M 43.44 13.72 45.85 15.45 3.40 29.23 66.13 0.21 23.65 26.90
+LayoutLMv3 [17] RoBERTa-B 138.4M 42.12 13.63 43.22 21.00 5.48 31.81 64.66 0.80 30.84 28.84
+DocLayout-YOLO [53] v10m 19.6M 43.71 48.71 72.83 42.67 35.40 51.44 64.64 9.54 57.54 47.38
+SwinDocSegmenter [4] Swin-L 223M 42.91 28.20 47.29 32.44 20.81 52.35 48.54 12.38 38.06 35.89
+GraphKD [5] R101 44.5M 39.03 16.18 39.92 22.82 14.31 37.61 44.43 5.71 23.86 27.10
+DOCX-Chain [50] - - 30.86 11.71 39.62 19.23 10.67 23.00 41.60 1.80 16.96 21.27
 Table 6: Component-level layout detection evaluation on OmniDocBench layout subset: mAP results by PDF page type.
-5.1 Component-specific Evaluation Results
 Model Type Model Language Table Frame Type Special Situation Overall
 EN ZH Mixed Full Omission Three Zero Merge Cell(+/-) Formula(+/-) Colorful (+/-) Rotate(+/-)
-OCR-based Models PaddleOCR 76.8 71.8 80.1 67.9 74.3 81.1 74.5 70.6/75.2 71.3/74.1 72.7/74.0 23.3/74.6 73.6
-RapidTable 80.0 83.2 91.2 83.0 79.7 83.4 78.4 77.1/85.4 76.7/83.9 77.6/84.9 25.2/83.7 82.5
-Expert VLMs StructEqTable 72.0 72.6 81.7 68.8 64.3 80.7 85.0 65.1/76.8 69.4/73.5 66.8/75.7 44.1/73.3 72.7
-GOT-OCR 72.2 75.5 85.4 73.1 72.7 78.2 75.7 65.0/80.2 64.3/77.3 70.8/76.9 8.5/76.3 74.9
-General VLMs Qwen2-VL-7B 70.2 70.7 82.4 70.2 62.8 74.5 80.3 60.8/76.5 63.8/72.6 71.4/70.8 20.0/72.1 71.0
-InternVL2-8B 70.9 71.5 77.4 69.5 69.2 74.8 75.8 58.7/78.4 62.4/73.6 68.2/73.1 20.4/72.6 71.5
+OCR-based Models PaddleOCR[23] 76.8 71.8 80.1 67.9 74.3 81.1 74.5 70.6/75.2 71.3/74.1 72.7/74.0 23.3/74.6 73.6
+RapidTable[37] 80.0 83.2 91.2 83.0 79.7 83.4 78.4 77.1/85.4 76.7/83.9 77.6/84.9 25.2/83.7 82.5
+Expert VLMs StructEqTable[55] 72.8 75.9 83.4 72.9 76.2 76.9 88.0 64.5/81.0 69.2/76.6 72.8/76.4 30.5/76.2 75.8
+GOT-OCR [45] 72.2 75.5 85.4 73.1 72.7 78.2 75.7 65.0/80.2 64.3/77.3 70.8/76.9 8.5/76.3 74.9
+General VLMs Qwen2-VL-7B [44] 70.2 70.7 82.4 70.2 62.8 74.5 80.3 60.8/76.5 63.8/72.6 71.4/70.8 20.0/72.1 71.0
+InternVL2-8B [8] 70.9 71.5 77.4 69.5 69.2 74.8 75.8 58.7/78.4 62.4/73.6 68.2/73.1 20.4/72.6 71.5
 Table 7: Component-level Table Recognition evaluation on OmniDocBench table subset. (+/-) means with/without special situation.
@@
 EN ZH Mixed White Single Multi Normal Rotate90 Rotate270 Horizontal
-Expert Vision Models PaddleOCR 0.071 0.055 0.118 0.060 0.038 0.085 0.060 0.015 0.285 0.021
-Tesseract OCR 0.179 0.553 0.553 0.453 0.463 0.394 0.448 0.369 0.979 0.982
-Surya 0.057 0.123 0.164 0.093 0.186 0.235 0.104 0.634 0.767 0.255
-GOT-OCR 0.041 0.112 0.135 0.092 0.052 0.155 0.091 0.562 0.966 0.097
-Mathpix 0.033 0.240 0.261 0.185 0.121 0.166 0.180 0.038 0.185 0.638
-Vision Language Models Qwen2-VL 0.072 0.274 0.286 0.234 0.155 0.148 0.223 0.273 0.721 0.067
-InternVL2 0.074 0.155 0.242 0.113 0.352 0.269 0.132 0.610 0.907 0.595
-GPT4o 0.020 0.224 0.125 0.167 0.140 0.220 0.168 0.115 0.718 0.132
+Expert Vision Models PaddleOCR [23] 0.071 0.055 0.118 0.060 0.038 0.085 0.060 0.015 0.285 0.021
+Tesseract OCR  [^5] https://github.com/tesseract-ocr/tesseract 0.179 0.553 0.553 0.453 0.463 0.394 0.448 0.369 0.979 0.982
+Surya  [^6] https://github.com/VikParuchuri/surya 0.057 0.123 0.164 0.093 0.186 0.235 0.104 0.634 0.767 0.255
+GOT-OCR [45] 0.041 0.112 0.135 0.092 0.052 0.155 0.091 0.562 0.966 0.097
+Mathpix 4 0.033 0.240 0.261 0.185 0.121 0.166 0.180 0.038 0.185 0.638
+Vision Language Models Qwen2-VL-72B [44] 0.072 0.274 0.286 0.234 0.155 0.148 0.223 0.273 0.721 0.067
+InternVL2-76B [8] 0.074 0.155 0.242 0.113 0.352 0.269 0.132 0.610 0.907 0.595
+GPT4o [2] 0.020 0.224 0.125 0.167 0.140 0.220 0.168 0.115 0.718 0.132
 Table 8: Component-level evaluation on OmniDocBench OCR subset: results grouped by text attributes using the edit distance metric.
 Models CDM ExpRate@CDM BLEU Norm Edit
-GOT-OCR 74.1 28.0 55.07 0.290
-Mathpix 86.6 2.8 66.56 0.322
-Pix2Tex 73.9 39.5 46.00 0.337
-UniMERNet-B 85.0 60.2 60.84 0.238
-GPT4o 86.8 65.5 45.17 0.282
-InternVL2 67.4 54.5 47.63 0.308
-Qwen2-VL 83.8 55.4 53.71 0.285
+GOT-OCR [45] 74.1 28.0 55.07 0.290
+Mathpix 4 86.6 2.8 66.56 0.322
+Pix2Tex  [^7] https://github.com/lukas-blecher/LaTeX-OCR 73.9 39.5 46.00 0.337
+UniMERNet-B [40] 85.0 60.2 60.84 0.238
+GPT4o [2] 86.8 65.5 45.17 0.282
+InternVL2-76B [8] 67.4 54.5 47.63 0.308
+Qwen2-VL-72B [44] 83.8 55.4 53.71 0.285
 Table 9: Component-level formula recognition evaluation on OmniDocBench formula subset.
-The OmniDocBench dataset features comprehensive and precise annotations, allowing for a fair and rigorous comparison of various document content extraction algorithms in real-world scenarios.
+4.3 Metric Calculation
+Pure Text. We calculate Normalized Edit Distance [21], averaging these metrics at the sample level to obtain the final scores.
+Tables. All tables are converted to HTML format before calculating the Tree-Edit-Distance-based Similarity (TEDS) [54] metric and Normalized Edit Distance.
+Formulas. Formulas are currently evaluated using the Character Detection Matching (CDM) metric [41], Normalized Edit Distance, and BLEU [33].
+Reading Order.
+Reading order is evaluated using the Normalized Edit Distance as metric.
+It only involves text components, with tables, images, and ignored components excluded from the final reading order calculation.
+5 Benchmarks
 Based on the distinct characteristics of these algorithms, we categorize document content extraction methods into three main classes:
-Pipeline Tools.
-These methods integrate layout detection and various content recognition tasks (such as OCR, table recognition, and formula recognition) into a document parsing pipeline for content extraction.
-Prominent examples include MinerU [36], Marker [30], and Mathpix  [^4] https://mathpix.com/.
-Expert VLMs. These are large multimodal models specifically trained for document parsing tasks. Representative models include GOT-OCR2.0 [39] and Nougat [5].
-General VLMs. These are general-purpose large multimodal models inherently capable of document parsing. Leading models in this category include GPT-4o [2], Qwen2-VL [38], and InternVL2 [6].
+  • Pipeline Tools: These methods integrate layout detection and various content recognition tasks (such as OCR, table recognition, and formula recognition) into a document parsing pipeline for content extraction.
+Prominent examples include MinerU [42] (v0.9.3), Marker [34] (v1.2.3), and Mathpix4.
+  • Expert VLMs: These are large multimodal models specifically trained for document parsing tasks. Representative models include GOT-OCR2.0 [45] and Nougat [7].
+  • General VLMs: These are general-purpose large multimodal models inherently capable of document parsing.
+Leading models in this category include GPT-4o [2], Qwen2-VL-72B [44], and InternVL2-76B [8].
 5.1 End-to-End Evaluation Results
-Utilizing the OmniDocBench dataset and our evaluation framework, we conducted end-to-end assessments of mainstream document parsing methods, evaluating their performance from input PDF images to the resultant document parsing outputs.
 Overall Evaluation Results.
-As illustrated in Table 2, pipeline tools specifically designed for document parsing, demonstrate superior performance across the board.
-MinerU and Mathpix achieved the best results for English and Chinese pages, respectively.
-In contrast, even the best general-purpose Vision Language Models (VLMs), GPT-4o, exhibits a performance gap compared to these specialized models, especially in Chinese.
-This trend is evident across sub-tasks like text recognition, formula recognition, and table recognition, where methods tailored for document parsing consistently outperform others.
-This advantage is largely due to the fine-tuning of these models on large datasets specific to document parsing tasks.
+As illustrated in Table 2, pipeline tools such as MinerU and Mathpix, demonstrate superior performance across sub-tasks like text recognition, formula recognition, and table recognition.
+Moreover, the general Vision Language Models (VLMs), Qwen2-VL, and GPT4o, also exhibit competitive performance.
+Almost all algorithms score higher on English than on Chinese pages.
 Performance Across Diverse Page Types.
 To gain deeper insights into model performance on diverse document types, we evaluated text recognition tasks across different page types.
-As shown in Table 3, an intriguing finding emerged: For commonly used data, such as academic papers and financial reports, pipeline tools perform well.
-However, for more specialized data like slides and handwritten notes, general VLMs demonstrate stronger generalization.
-The reason is clear: Pipeline tools and expert VLMs are relatively more constrained by the range of training data, whereas general VLMs having been trained on a wide variety of samples, maintained excellent recognition performance even in traditionally challenging long-tail scenarios, underscoring the value of VLMs.
-Performance on Pages with Specific Attributes.
-For documents in OmniDocBench with attributes such as fuzzy scans, watermarks, and colorful backgrounds, our evaluation results are presented in Table 4.
-In these scenarios, the VLMs InternVL2 and Qwen2-VL exhibit the strongest resistance to interference, achieving the best accuracy and robustness.
-MinerU also performs commendably.
-Performance on Different Column Layout Types.
-OmniDocBench annotates page attributes such as column layout type, which is crucial for analyzing model performance in reading order.
-As depicted in Table 5, all models experience a noticeable decline in reading order accuracy when dealing with complex layouts.
-MinerU and Mathpix excels in reading order across various column layouts, demonstrating robust performance across different page types.
-From these end-to-end evaluations, it is evident that pipeline tools like MinerU and Mathpix, specifically designed for document parsing, achieve the best overall performance.
-However, in terms of versatility and scalability, VLMs offer a distinct advantage over pipeline tools.
-Fine-tuning a general large model like Qwen2-VL with specialized data could yield models even more adept at document parsing, indicating a promising direction for future research in multimodal approaches.
-The OmniDocBench dataset provides comprehensive annotations for document parsing, including layout detection, text boxes and content, formula boxes and content, and table boxes and content.
-These detailed annotations enable the evaluation of current state-of-the-art (SOTA) methods across various document types, allowing us to analyze their performance in diverse scenarios.
-Additionally, these results can be used to assemble enhanced pipeline tools for document parsing tasks.
-5.3 Single Algorithm Evaluation Results
+Intriguingly, as shown in Table 3, pipeline tools perform well for commonly used data, such as academic papers and financial reports.
+Meanwhile, for more specialized data, such as slides and handwritten notes, general VLMs demonstrate stronger generalization.
+Notably, most VLMs fail to recognize when dealing with the Newspapers, while pipeline tools achieve significantly better performance.
+Performance on Pages with Visual Degradations.
+In Table 4, we further analyze performance on pages containing common document-specific challenges, including fuzzy scans, watermarks, and colorful backgrounds.
+VLMs like InternVL2 and Qwen2-VL exhibit higher robustness in these scenarios despite visual noise.
+Among pipeline tools, MinerU remains competitive due to its strong layout segmentation and preprocessing capabilities.
+Performance on Different Layout Types.
+Page layout is a critical factor in document understanding, especially for tasks involving reading order.
+OmniDocBench annotates layout attributes such as single-column, multi-column, and complex custom formats.
+Across all models, we observe a clear drop in accuracy on multi-column and complex layouts.
+MinerU shows the most consistent reading order prediction, though its performance dips on handwritten single-column pages due to recognition noise.
+Discussion on End-to-End Results.
+1) While general VLMs often lag behind specialized pipelines and expert models on standard documents (e.g., academic papers), they generalize better to unconventional formats (e.g., notes) and perform more robustly under degraded conditions (e.g., fuzzy scans).
+This is largely due to their broader training data, enabling better handling of long-tail scenarios compared to models trained on narrow domains.
+2) VLMs, however, struggle with high-density documents like newspapers due to limitations in input resolution and token length.
+In contrast, pipeline tools leverage layout-based segmentation to process components individually, maintaining accuracy in complex layouts.
+Enhancing VLMs with layout-aware designs and domain-specific fine-tuning offers a promising path forward.
+OmniDocBench facilitates this by providing detailed annotations for layout, text, formulas, and tables, enabling comprehensive benchmarking and modular tool development for diverse document parsing tasks.
+5.2 Single Task Evaluation Results
 Layout Detection Results.
@@
 This superiority is a key factor in MinerU’s integration of DocLayout-YOLO, contributing to its outstanding overall performance.
-The table also reveals that, aside from DocLayout-YOLO, other methods perform well on books and academic literature but are less effective on other document types, primarily due to a lack of pre-training on diverse documents.
+Other methods perform well on books and academic literature but struggle with more diverse formats due to limited training data.
 Table Recognition Results.
-Table recognition results evaluated by Tree-Edit-Distance-based Similarity (TEDS) metric are presented in Table 7.
-We evaluate table recognition models across three dimensions on our OmniDocBench table subset: language diversity, table frame types, and special situations.
+In Table 7, We evaluate table recognition models across three dimensions on our OmniDocBench table subset: language diversity, table frame types, and special situations.
 Among all models, OCR-based models demonstrate superior overall performance, with RapidTable achieving the highest scores in language diversity and maintaining stable performance across different frame types.
-Expert VLMs show competitive results in specific scenarios, with StructEqTable [48] excelling in no frame tables and showing better rotation robustness.
+Expert VLMs show competitive results in specific scenarios, with StructEqTable [55] excelling in no-frame tables and showing better rotation robustness.
 General VLMs (Qwen2-VL-7B and InternVL2-8B) exhibit relatively lower but consistent performance, suggesting that while general-purpose VLMs have made progress in table understanding, they still lag behind specialized solutions.
 Text Recognition Results.
-In the traditional OCR task, Table 8 shows that PaddleOCR leads the field, surpassing other models significantly, with GOT also performing relatively well.
-Selecting these two methods for the OCR module is a prudent choice.
+Table 8 compares OCR tools across languages, backgrounds, and rotations using Edit Distance.
+PaddleOCR outperforms all competitors, followed by GOT-OCR and Mathpix.
+General VLMs struggle to handle text rotation or mixed-language scenarios.
 Formula Recognition Results.
-For formula recognition, the CDM metric provides a clear comparison in Table 9.
+Table 9 presents results on formula parsing, using CDM, BLEU, and normalized Edit Distance.
 GPT-4o, Mathpix, and UniMERNet achieve results of 86.8%, 86.6%, and 85.0%, respectively.
@@
 The results show that table rotation significantly impacts the accuracy of all models.
-Pipeline Tools perform well on more challenging tables, but colored backgrounds can affect recognition accuracy.
+Pipeline Tools’ performance would not be affected by more challenging tables (e.g., merge cell), but colored backgrounds can affect recognition accuracy.
 Several Vision Language Models (VLMs) tend to perform worse on tables with merged cells, but colored backgrounds do not significantly impact table recognition accuracy.
@@
 Some models, such as MinerU and Marker, experience a further decrease in accuracy when recognizing mixed Chinese and English content.
-Complex background colors significantly affect the recognition accuracy of pipeline tools, but they have little impact on VLMs.
+The main reason is that minerU’s text recognition module is PaddleOCR model.
+According to the performance of the PaddleOCR model in text recognition module, its accuracy will decline in the case of mixed language.
+Moreover, complex background colors significantly affect the recognition accuracy of pipeline tools, but it has only little impact on accuracy for VLMs.
 Model Type Model Language Table Frame Type Special Situation
 EN ZH Mixed Full Omission Three Zero Merge Cell(+/-) Formula(+/-) Colorful (+/-) Rotate(+/-)
-Pipeline Tools MinerU 75.7 59.9 79.6 60.0 72.8 70.1 60.4 64.1/66.0 66.7/65.0 59.8/68.1 2.9/66.4
-Marker 52.5 43.0 44.2 41.8 55.3 47.1 52.4 43.8/47.0 42.9/46.6 44.3/46.7 6.3/46.6
-Mathpix 76.1 64.3 71.9 68.3 79.3 67.0 25.8 71.2/66.4 69.8/67.6 60.5/71.8 20.7/68.8
-Expert Vision Models GOT-OCR 51.9 47.0 49.4 46.2 49.3 51.6 47.2 46.5/49.7 46.4/49.1 40.2/52.7 0.0/49.4
-Nougat 36.5 0.4 0.0 6.3 3.6 22.2 0.0 15.1/9.1 21.2/8.9 2.8/15.3 0.0/11.4
-Vision Language Models GPT4o 71.8 58.8 57.9 63.3 69.5 61.9 31.8 57.5/65.5 61.6/62.9 62.0/63.0 14.5/63.5
-Qwen2-VL 57.4 62.9 72.7 70.7 64.1 48.3 57.6 49.4/68.2 48.5/64.7 63.5/60.7 41.6/61.9
-InterVL2 61.5 59.3 65.9 59.7 66.5 58.7 56.2 49.6/65.9 54.4/61.6 59.4/60.6 7.3/61.1
+Pipeline Tools MinerU 75.1 59.3 79.1 59.4 71.6 69.7 60.0 63.6/65.3 66.0/64.4 59.2/67.5 3.0/65.8
+Marker 64.9 47.3 49.8 44.5 61.8 59.0 63.6 52.6/52.7 53.2/52.5 48.0/54.9 35.5/52.9
+Mathpix 75.4 63.2 71.3 67.4 77.3 66.3 25.5 70.3/65.4 68.7/66.7 59.7/70.8 19.2/67.9
+Expert Vision Models GOT-OCR 51.7 46.2 49.0 45.5 48.3 51.3 46.2 46.0/48.9 45.7/48.4 39.8/51.9 0.0/48.7
+Nougat 36.2 0.3 0.0 6.1 3.5 22.1 0.0 15.0/8.9 21/8.7 2.6/15.2 0.0/11.2
+Vision Language Models GPT4o 71.1 58.0 57.3 62.5 68.7 61.3 31.2 56.8/64.7 60.8/62.2 61.4/62.2 14.2/62.7
+Qwen2-VL-72B 73.2 75.1 76.1 72.0 79.0 77.5 63.2 67.9/78.1 71.6/75.3 77.9/72.9 42.7/75.1
+InterVL2-76B 60.9 58.5 65.4 58.8 65.3 58.3 55.6 49.0/65.1 53.3/60.9 58.8/59.8 6.9/60.3
 Table S1: End-to-End Table TEDS Result grouped by Table Attributes
@@
 EN ZH Mixed White Single Multi
-Pipeline Tools MinerU 0.123 0.206 0.742 0.163 0.147 0.513
-Marker 0.267 0.389 0.499 0.339 0.389 0.497
-Mathpix 0.173 0.774 0.538 0.675 0.554 0.570
+Pipeline Tools MinerU 0.124 0.234 0.742 0.188 0.15 0.514
+Marker 0.163 0.379 0.747 0.303 0.396 0.594
+Mathpix 0.175 0.793 0.538 0.698 0.587 0.583
 Expert Vision Models GOT-OCR 0.251 0.763 0.266 0.669 0.595 0.440
@@
 Vision Language Models GPT4o 0.170 0.647 0.322 0.536 0.423 0.406
-Qwen2-VL 0.337 0.575 0.310 0.537 0.400 0.233
-InternVL2 0.418 0.606 0.251 0.589 0.366 0.221
+Qwen2-VL-72B 0.128 0.582 0.209 0.494 0.388 0.217
+InternVL2-76B 0.418 0.606 0.251 0.589 0.366 0.221
 Table S2: End-to-End Text Normalized Edit Distance results grouped by Text Attributes. “Mixed” represents a mixture of Chinese and English, “Single” and “Multi” represent single color and multi color.
-II Dataset Statistics and Visualization
-OmniDocBench contains 981 pages, including 9 types of PDF pages, 4 types of layouts, and 3 types of languages.
-Some pages also include special conditions, such as watermarks.
-Table S3 and Figure S1 show the number of pages with each page attribute.
-Figures S3, S4, S5 and S6 are examples of PDF pages with different PDF types, Layout Types, and Special Issues.
-Table S6 and Figure S2 show all annotation categories included in OmniDocBench.
-All of them are annotated by bounding boxes.
-There are 15 types of block-level annotations and 4 types of span-level annotations, with span-level annotations nested within the block-level ones.
-In addition, there are 3 types of annotations marked as page interference information (No.20-22), whose bounding boxes are used to mask the specific regions of the PDF pages to avoid affecting the evaluation results.
-The recognition annotations are also provided for each annotation category except for Figures.
-Formulas is written in LaTeX format and Table is annotated in both HTML and LaTeX formats.
-Others are annotated in plain text.
-Furthermore, the Text Attributes are also annotated for each block-level category that contains text.
-There are 3 types of Text Attributes that might influent OCR accuracy: Language, Text Background Color, and Text Rotation.
-Table S5 shows the statistics of annotations with specific text attributes.
-There are 23,010 block-level annotations are labeled with text attributes.
-Tables are also annotated with Table Attributes.
-There are 6 types of Table Attributes that might influent the Table Recognition accuracy: Language, Table Frame Type, Merge Cell, Colorful Background, Contain Formula, and Rotation.
-Table S5 shows the numbers of annotations with specific table attributes.
-Figures S7 and S8 are the examples of Tables with different Frames and Special Issues.
-III Model Results Visualization
-Figures S10, S11, S12, S13, S14, S9, S15, S16 and S17 show the examples of Good model outputs and Bad model outputs of Document Parsing among different PDF types.
-As it shown, different models exhibit varying performance across different PDF types.
-For example, MinerU detects all handwritten notes as figures, resulting in very low recognition accuracy in Notes.
-Marker and InternVL2 experience missed detections, leading to lower scores.
-InternVL2 and Qwen2-VL, in specific PDF types (such as slides or financial reports), tend to merge multi-column text.
-Figures S20, S18 and S19 show the examples of Good model outputs and Bad model outputs under special issues of the PDF pages.
-It shows that Marker tends to generate typos when the PDF pages are fuzzy scanned or with watermarks, while GOT-OCR fails to recognize content on pages with colored backgrounds.
-MinerU performs well under special situations, while Mathpix occasionally generates typos.
-Figures S21, S22, S23 and S24 show examples of Good model outputs and Bad model outputs for PDF pages with different layouts.
-MinerU has a low reading order score for single-column layouts primarily because most notes are single-column, and MinerU performs poorly in recognizing Notes, leading to a low reading order score accordingly.
-InternVL2 scores high in Single-Column layouts but scores poorly on Double-Column and Three-Column layouts.
-It is mainly due to frequent missed content recognition and errors in reading order judgment in multi-column layouts pages.
-MinerU’s reading order and recognition accuracy decrease with complex layouts, primarily because it incorrectly merges multiple columns during recognition.
-Figures S27 and S28 show the model’s recognition ability under special issues of text.
-In text recognition with complex background colors, Marker may produce errors or miss content, whereas Qwen2-VL still performs well.
-Most models fail to recognize text when it is rotated 270 degrees.
-Some vision language models generate hallucinated information based on the content they can recognize.
-Figures S29, S30, S31 and S32 show the examples of good and bad model results for tables with different attributes.
-For three-line tables, RapidTable demonstrates a good performance with accurate structure recognition, while PaddleOCR shows limitations by missing the last column in its outputs.
-Interestingly, in tables without frames, PaddleOCR performs well with accurate table predictions, while Qwen2-VL-7B exhibits errors in the last two columns.
-This indicates that the presence or absence of table frames can significantly impact different models’ performance in different ways.
-Rotated tables prove to be particularly challenging, with most models, including GOT-OCR, failing to recognize the table structure.
-However, StructEqTable shows promising results by correctly identifying most of the table content, though with a few detail errors.
-For tables containing formula, Qwen2-VL-7B shows more accurate table structure recognition compared to InternVL2-8B.
-IV Model Settings
-For pipeline tools such as MinerU, Marker, and Mathpix, default settings are used for evaluation.
-Specifically, MinerU with Version 0.9.3555https://github.com/opendatalab/MinerU/releases/tag/magic_pdf-0.9.3-released is employed.
-For Marker, Version 0.2.17666https://github.com/VikParuchuri/marker/releases/tag/v0.2.17 is evaluated.
-For Nougat, we utilize its 0.1.0-base model (350M).
-For GOT-OCR, we employ its format OCR mode to output structured data.
-For general VLMs, we used the GPT4o, Qwen2-VL-72B, and InternVL2-Llama3-76B by setting the do_sample===False to ensure the reproducibility.
 Category Attribute Name Count
@@
 Table S6: Annotation Explanations and Statistics.
+II Dataset Statistics and Visualization
+OmniDocBench contains 981 pages, including 9 types of PDF pages, 4 types of layouts, 3 types of languages, and 3 special issues in visual degradations (e.g., watermarks).
+Table S3 and Figure S1 show the number of pages with each page attribute.
+Figures S5, S6, S7 and S8 are examples of PDF pages with different PDF types, Layout Types, and Special Issues.
+Table S6 and Figure S2 show all annotation categories included in OmniDocBench.
+All of them are annotated by bounding boxes.
+There are 15 types of block-level annotations and 4 types of span-level annotations, with span-level annotations nested within the block-level ones.
+In addition, there are 3 types of annotations marked as page interference information (No.20-22), whose bounding boxes are used to mask the specific regions of the PDF pages to avoid affecting the evaluation results.
+The recognition annotations are also provided for each annotation category except for Figures.
+Formulas is written in LaTeX format and Table is annotated in both HTML and LaTeX formats.
+Others are annotated in plain text.
+Furthermore, the Text Attributes are also annotated for each block-level category that contains text.
+There are 3 types of Text Attributes that might influent OCR accuracy: Language, Text Background Color, and Text Rotation.
+Table S5 shows the statistics of annotations with specific text attributes.
+There are 23,010 block-level annotations are labeled with text attributes.
+Tables are also annotated with Table Attributes.
+There are 6 types of Table Attributes that might influent the Table Recognition accuracy: Language, Table Frame Type, Merge Cell, Colorful Background, Contain Formula, and Rotation.
+Table S5 shows the numbers of annotations with specific table attributes.
+Figures S9 and S10 are the examples of Tables with different Frames and Special Issues.
+III Discussion on Model Predictions
+Conclusion Combining scattered results from tasks and sub-attributes, it can be concluded that pipeline tools and expert models have better performance on common data like academic papers and challenging cases such as tables with merged cells compared to VLMs.
+However, VLMs demonstrate stronger generalization on uncommon PDF types like slides and exam papers, and they show greater robustness in special page situations, such as fuzzy scans.
+The low accuracy of VLMs is mainly due to:1) Missing Content in dense pages(Figure S11); 2) Hallucinations in hard-to-recognize pages(Figure S30).
+The low accuracy of Pipeline tools mainly due to: 1) Lower robustness in special page situations, e.g., watermark(Figure S21); 2) Weak generalization on uncommon PDF types, e.g., handwriting notes(Figure S16).
+Figures S12, S13, S14, S15, S16, S11, S17, S18 and S19 show the examples of Good model outputs and Bad model outputs of Document Parsing among different PDF types.
+As it shown, different models exhibit varying performance across different PDF types.
+For example, MinerU detects all handwritten notes as figures, resulting in very low recognition accuracy in Notes.
+Marker and InternVL2 experience missed detections, leading to lower scores.
+InternVL2 and Qwen2-VL, in specific PDF types (such as slides or financial reports), tend to merge multi-column text.
+Figures S22, S20 and S21 show the examples of Good model outputs and Bad model outputs under special issues of the PDF pages.
+It shows that Marker tends to generate typos when the PDF pages are fuzzy scanned or with watermarks, while GOT-OCR fails to recognize content on pages with colored backgrounds.
+MinerU performs well under special situations, while Mathpix occasionally generates typos.
+Figures S23, S24, S25 and S26 show examples of Good model outputs and Bad model outputs for PDF pages with different layouts.
+MinerU has a low reading order score for single-column layouts primarily because most notes are single-column, and MinerU performs poorly in recognizing Notes, leading to a low reading order score accordingly.
+InternVL2 scores high in Single-Column layouts but scores poorly on Double-Column and Three-Column layouts.
+It is mainly due to frequent missed content recognition and errors in reading order judgment in multi-column layouts pages.
+MinerU’s reading order and recognition accuracy decrease with complex layouts, primarily because it incorrectly merges multiple columns during recognition.
+Figures S29 and S30 show the model’s recognition ability under special issues of text.
+In text recognition with complex background colors, Marker may produce errors or miss content, whereas Qwen2-VL still performs well.
+Most models fail to recognize text when it is rotated 270 degrees.
+Some vision language models generate hallucinated information based on the content they can recognize.
+Figures S31, S32, S33 and S34 show the examples of good and bad model results for tables with different attributes.
+For three-line tables, RapidTable demonstrates a good performance with accurate structure recognition, while PaddleOCR shows limitations by missing the last column in its outputs.
+Interestingly, in tables without frames, PaddleOCR performs well with accurate table predictions, while Qwen2-VL-7B exhibits errors in the last two columns.
+This indicates that the presence or absence of table frames can significantly impact different models’ performance in different ways.
+Rotated tables prove to be particularly challenging, with most models, including GOT-OCR, failing to recognize the table structure.
+However, StructEqTable shows promising results by correctly identifying most of the table content, though with a few detail errors.
+For tables containing formula, Qwen2-VL-7B shows more accurate table structure recognition compared to InternVL2-8B.
+IV Model Settings
+For pipeline tools such as MinerU, Marker, and Mathpix, default settings are used for evaluation.
+Specifically, MinerU with Version 0.9.3888https://github.com/opendatalab/MinerU/releases/tag/magic_pdf-0.9.3-released is employed.
+For Marker, Version 1.2.3999https://github.com/VikParuchuri/marker/releases/tag/v1.2.3 is evaluated.
+For Nougat, we utilize its 0.1.0-base model (350M).
+For GOT-OCR, we employ its format OCR mode to output structured data.
+For general VLMs, we used the GPT4o, Qwen2-VL-72B, and InternVL2-Llama3-76B by setting the do_sample===False to ensure the reproducibility.
+After testing the different setting of max_token, the best setting is chosen for each VLMs.
+Specifically, max_token===32000 is set for Qwen2-VL-72B, and max_token===4096 is set for InternVL2-Llama3-76B.
+For GPT-4o, the default setting is used.
+V More Details on Methods
+Ignore handling.
+The purpose of this process is to avoid fluctuations in accuracy caused by the lack of uniformity in the output standards among document parsing algorithm.
+(1) Some algorithm (e.g., GPT-OCR, Qwen2-VL) tends to remove headers and footers, while others (e.g., GPT4o) prefers to retain them ( Figure S3).
+(2) Moreover, the reading order mismatch cause by captions and footnotes is also considered.
+For example, Nougat would put the image captions in the end of the page content( Figure S4), while others tend to put the image captions in human reading order.
+Ignore handling is to minimize the impact of varying standards of document parsing on evaluation.
+Our evaluation dataset aims to more fairly assess the parsing accuracy of various algorithms, and these trivial issues regarding standards are not within our scope of consideration.
 Figure S1: The Data Proportion of Pages for each Attribute in OmniDocBench.
 Figure S2: The Visualization of vary Annotations in OmniDocBench.
+Figure S3: The Vary Standards in parsing Header, Footers, and so on.
+Figure S4: The Vary Standards in parsing Captions.
 Figure S5: The Examples of Academic Papers, Books, Textbooks, Notes, and Magazines in OmniDocBench.